### v1

# A template for the arxiv style

Abstract
We consider the task of identifying the causal parents of a target variable among a set of candidate variables from observational data. Our main assumption is that the candidate variables are observed in different environments which may, for example, correspond to different settings of a machine or different time intervals in a dynamical process. Under certain assumptions different environments can be regarded as interventions on the observed system. We assume a linear relationship between target and covariates, which can be different in each environment with the only restriction that the causal structure is invariant across environments. This is an extension of the ICP (Invariant Causal Prediction) principle by Peters et al. (2016), who assumed a fixed linear relationship across all environments. Within our proposed setting we provide sufficient conditions for identifiability of the causal parents and introduce a practical method called LoLICaP (Locally Linear Invariant Causal Prediction), which is based on a hypothesis test for parent identification using a ratio of minimum and maximum statistics. We then show in a simplified setting that the statistical power of LoLICaP converges exponentially fast in the sample size, and finally we analyze the behavior of LoLICaP experimentally in more general settings.
1 Introduction
We consider the problem of identifying the causal parents of a target variable, among a set of candidate variables, based only on observational data. The main assumption we make is that data is observed in different environmental scenarios. An emblematic example is that of machine diagnostics, where we are monitoring several connected components of the machine. Different environments correspond, for example, to machines in different locations, different settings of a machine, or data collected in different points in time. If the system behaves differently across environments we talk about heterogeneous environments, and these can be interpreted as accidental interventions on the system. This means that, under certain modeling assumptions, one can leverage the information provided by those accidental interventions to perform causal discovery tasks. One particular way of using this information is based on the idea that the performance of any prediction model for the target variable should be invariant under interventions (in different environments) if and only if this prediction model uses exclusively the causal parents of the target: For example, a model that uses the color green to predict that an image contains a cow may work well if we only have cow pictures from grassy areas. The green background is, however, only a correlated variable and not a causal parent, and the performance of such a model should significantly drop if we suddenly present it with cow images with a sandy background. This principle of invariant causal prediction (ICP) was proposed by Peters et al. (2016), who assumed that the causal relationship between the target and its parents follows a linear model which is fixed across all environments. We extend upon that work by relaxing the global linearity assumption to a locally linearity assumption, meaning that each environment is equipped with its own linear model. Relaxing the global model to local models carries a couple of interesting implications that we address in this paper. In particular, locally linear models extend the scope of what might be considered heterogeneous environments. Heterogeneity as viewed in Peters et al. (2016) is always introduced by changes in the feature distributions. Considering locally linear models, heterogeneity can also be induced by parameter changes in the local models, which can be seen as changes in the causal mechanism between different environments. The main contributions of this paper are as follows:
  • We extend the work of Peters et al. (2016) to a locally linear model class.
  • We provide sufficient conditions under which the causal parents of a target variable are identifiable within the model class under investigation.
  • Given a finite amount of data, we introduce LoLICaP (Locally Linear Invariant Causal Prediction), a method suitable to decide if a given set of potential causal parents is a plausible causal explanation of a target variable.
  • Under certain normality assumptions we show that the power of LoLICaP converges exponentially fast, in the sample size, towards one.
  • We provide experimental evidence that LoLICaP is applicable in a broad scope of settings.
Organization:
In Section 2 we contextualize our contributions by discussing related work. In Section 3 we introduce the setting, our assumptions and formalize inference goals. This section features, in addition, concrete examples illustrating the meaningfulness of the modeling assumptions. In Section 4 we present a meta-approach and give a characterization of its potential in terms of identifying causal relations. Motivated by that, Section 5 presents a concrete approach to detect causal parents with limited amount of data. Section 6 illustrates the performance of LoLICaP in a variety of settings. We conclude the paper in Section 7 with a discussion and an outlook on future work.
2 Related work
The idea of relying on heterogeneous environments for causal discovery is not new, and was instrumental for the work in Peters et al. (2016). This setup was also extended to sequential data in Pfister et al. (2019), while Heinze-Deml et al. (2018) investigate non-linear models by using conditional independence tests. Some papers relax the global linearity assumption for local models with the help of structural assumptions between the local models: Christiansen and Peters (2020) and Zhou et al. (2022) assume that the local structural parameters, i.e. the parameters of the assumed linear relationship between target and its parents, are related through a common (unobserved) variable. Huang et al. (2019) considers a temporal setting, where the local structural parameters change according to an auto-regressive model. The same authors extent this to a constraint based method under the assumption that the changes in the parameters happen independently (Huang et al., 2020). In contrast to these works we consider a setting where the structural parameters can be radically different for the various environments.
The concept of causal invariance is also carried to the context of out-of-domain generalization for machine learning models (Arjovsky et al., 2019; Oberst et al., 2021; Wald et al., 2021). The main underlying idea is that causal features, in contrast to merely correlated features, should be useful in all domains. In other words, the predictive power of causal features should be invariant in all domains. Differences in the approaches in that literature often stem from which parts of the model are assumed to be invariant across environments. While the goal of out-of-domain generalization is different than our goal, which is causal discovery, many of the concepts and models resemble each other.
The following papers aggregate the information of different environments, without needing the heterogeneity aspect, while typically assuming that the causal structure is acyclic, and thus representable by a directed acyclic graph (DAG). Osman et al. (2023) propose a method that finds causal structures and detects interventions using a minimum description length score on the causal factorization, which they define jointly over the different environments. They assume non-linear relationships, identifiability of the DAG up the Markov equivalence class and a low noise condition. Chen et al. (2021) assume a linear structural equation model (SEM), such that at least in some environments the corresponding DAG is identifiable from data. To find a solution of the DAG they propose a penalized maximum likelihood estimator which is optimized jointly over all environments. Shimizu (2012) proposes an extension of LiNGAM (Shimizu et al., 2006) to the multiple environment case. LiNGAM is a method that can find DAGs under a linearity and non-Gaussianity assumption.
Finally, Mooij et al. (2020) developed the joint causal inference (JCI) framework, where environments (in the paper called contexts) are directly encoded as part of the structural causal model (SCM). Many existing methods, in particular also the work of Peters et al. (2016), can be stated in that framework, and many methods for single environments, e.g. fast causal inference by Spirtes (2001), can be used within the JCI framework.
3 Setting
In this work we consider a scenario where there are two observable quantities of interest: a target, denoted by $Y$, and a set of covariates $X\vcentcolon=(X_{1},\cdots,X_{D})$ for $D\in\mathbb{N}$. Our overarching goal is to identify which of the $D$ covariates are the causal parents of the target $Y$. What we precisely mean with a causal parent is detailed in the next subsection. To specify the relation we assume between target and covariates, we introduce the notion of an environment. Specifically, suppose we have $E\in\mathbb{N}$ different environments and in each environment we receive $n^{e}$ observations, so that $Y^{e}\in\mathbb{R}^{n^{e}}$ and $X^{e}\in\mathbb{R}^{n^{e}\times D}$ are respectively all the target and covariate observations. Our main assumption is that for each $e\in[E]$ there exists a column vector $\beta^{e}\in\mathbb{R}^{D}$ such that
$Y^{e}=X^{e}\beta^{e}+\varepsilon^{e},$ (1)
where $\varepsilon^{e}\in\mathbb{R}^{n^{e}}$ is a zero-mean random noise. The specific independence assumptions on $\varepsilon^{e}$ are made explicit below. In this work we consider the assumed relationship $Y^{e}=X^{e}\beta^{e}+\varepsilon^{e}$ to be a structural causal model (SCM) in the sense of Pearl (2016), and thus we consider the set $S^{e,*}\vcentcolon=\{d\in[D]\mid\beta^{e}_{d}\neq 0\}\subseteq[D]$ as the true causal parents of $Y^{e}$. This implies that $Y^{e}$ can have different causal parents in each environment, and we consider a variable with index $d\in[D]$ a causal parent of $Y$ if $d\in S^{e,*}$ for at least one $e\in[E]$. Formally we define the set of causal parents of $Y$ as $S^{*}\vcentcolon=\bigcup_{e\in[E]}S^{e,*}$. Our inference goal of finding the causal parents of $Y$ is then equivalent to finding $S^{*}$.
3.1 Notation
This section introduces notation we will use throughout the paper. A graph $G=([D],\mathcal{E})$ is a tuple where indices in $[D]$ represent nodes and $\mathcal{E}\subseteq[D]^{2}$ represent directed edges between nodes, with the assumption that $(d,d)\not\in\mathcal{E}$ for any $d\in[D]$. If $(d_{1},d_{2})\in\mathcal{E}$ we call $d_{1}$ a parent of $d_{2}$ and we call $d_{2}$ a child of $d_{1}$. We call $d_{1}$ an ancestor of $d_{k}$ and we call $d_{k}$ a descendant of $d_{1}$ if there exists a sequence $d_{1},\cdots,d_{k}$ such that $(d_{i},d_{i+1})\in\mathcal{E}$ for $1\leq i<k$. If $d_{k}$ is not a descendant of $d_{1}$ we call $d_{k}$ a non-descendant of $d_{1}$. A node without any child is called a sink node. Given a node $d\in[D]$ we respectively define $\textbf{PA}(d)$, $\textbf{AN}(d)$, $\textbf{DE}(d)$, $\textbf{NDE}(d)$ as the set of all parents, ancestors, descendants and non-descendants of $d$.
Some further notation we use: With $X^{e}_{d,i}$ we indicate the entry of $X^{e}$ in the $d$-th row and $i$-th column, which corresponds to the $d$-th covariate of observation $i$. Similarly $\beta^{e}_{d}$ indicates the $d$-th entry of $\beta^{e}$. For $S\subseteq[D]$ we write $X^{e}_{S}\in\mathbb{R}^{n^{e}\times|S|}$ to indicate the sub-matrix of $X^{e}$ with columns given by $S$. In our notation the superscript $e$ always indicates the corresponding environment. We use $\mathbf{E}[\cdot]$, $\mathbf{V}[\cdot]$ and $\mathbf{C}[\cdot,\cdot]$ respectively as the expectation, variance and covariance operator. Finally we set $(\boldsymbol{X},\boldsymbol{Y})\vcentcolon=\{(X^{e},Y^{e})\}_{e\in[E]}$.
3.2 Data generation mechanism
Figure 1: The above graph shows the causal relations we assume, shown by the solid arrows. The dotted arrows may exist, but are not assumed. When we generate data from this type of model, we assume that we have access to $E$ different environments $e$ with structural parameter $\beta^{e}$ (outer box), and for each environment $e$ we observe $n^{e}$ samples according to Equation (1) (inner box).
While the previous subsections outline the main idea of our setting, we formalize it in the following assumption:
Assumption 1.
There exists a subset $S^{*}\subseteq[D]$, a zero mean distribution $F^{*}$ and a collection of coefficients $\beta^{e}\in\mathbb{R}^{D}$ for $e\in[E]$, such that $\beta^{e}_{d}=0$ for all $d\not\in S^{*}$ and $e\in[E]$. Furthermore, for all $e\in[E]$ we have that
$Y^{e}=X^{e}\beta^{e}+\varepsilon^{e}\ \ \text{ \ with: \ }$:
$\varepsilon^{e}_{i}\sim F^{*}\ \ \text{\ for all $i\in[n^{e}]$}$ (2)
$\varepsilon^{e}_{i}\perp\!\!\!\perp\varepsilon^{e}_{j}\ \ \text{\ for all $i\neq j$\ }$
$\varepsilon^{e}_{i}\perp\!\!\!\perp X^{e}_{S^{*},i}\ \ \text{\ for all $i\in[n^{e}]$}.$ (3)
[…]
At first, our setting may seem like a sparse regression task, but we emphasize that this model is considered to be a structural causal model, and the parameters $\beta^{e}$ have the interpretation of structural, rather than regression coefficients. We refer to Section 3.8.1 in Pearl (2016) for further discussion on this distinction. A particular consequence is that in condition (3) we require the unobserved noise $\varepsilon^{e}$ to be only independent of the causal parents $X^{e}_{S^{*}}$, but not necessarily of other variables. This becomes important when we want to find causal relations, instead of only statistical associations. Consider, for example, the simple causal structure of Figure 2. Here $X^{e}_{1}$ is a causal parent of $Y^{e}$ and independent of $\varepsilon^{e}$. However, $X^{e}_{2}$ is a descendant of $Y^{e}$ and thus not necessarily independent of $\varepsilon^{e}$.
We also draw special attention to condition (2), i.e., the assumption that $\varepsilon^{e}_{i}\sim F^{*}$ for all $e\in[E]$ and $i\in[n^{e}]$. This is arguably one of the stronger assumptions in our setting, and could be relaxed as discussed in Section 7. The main purpose of this assumption is, however, to ensure the environment $e$ does not have any influence on the unobserved variables. Considering the environment again as an intervention it becomes clear that some form of this assumption is necessary, as otherwise we have an unobserved confounding effect of the intervention.
It is also important to highlight that we do not make any independence assumptions across the different environments. The assumption that the distribution $F^{*}$ has zero mean is not crucial. This can always be achieved by adding an extra column to $X^{e}$, which is deterministically one, while the corresponding structural parameter in $\beta^{e}$ is then the negative mean of $F^{*}$.
Figure 2: In this simple model the variable $X^{e}_{2}$ is not necessarily independent of the noise $\varepsilon^{e}$.
Finally, we mention at several stages that the different environments should be heterogeneous, i.e., they should be different from each other in some aspect. In view of Assumption 1, heterogeneity corresponds either to different distributions of the covariates $X^{e}$, to different parameters $\beta^{e}$, or both. Within the context of our paper heterogeneity of environments is needed to identify causal parents. In Section 4.2, however, we provide sufficient conditions that ensure that the causal parents are identifiable for almost all possible values of $(\beta^{e})_{e\in[E]}$.
3.2.1 Examples
We now describe two settings in which Assumption 1 is sensible, so that we could apply LoLICaP, which we describe later in the paper.
Finding causal parents of a disease.
Assume we want to find causes for a certain disease. We collect data from different countries, which serve as the environments. We then collect data from plausible risk factors for the disease, where we have to make sure that we include all factors, which are heterogeneous across environments, such that the unexplained variation of the disease is roughly of the same distribution in all environments. Heterogeneous causal parents can be for example the diet, stress factors such as working conditions, lifestyle, genetic variations etc. In this setting LoLICaP is, in principle, able to identify the true causes of the disease amongst all plausible risk factors.
Finding causes with temporal data and mechanism shift.
We are collecting dynamical data from a machine and observe that a target variable $Y$ starts to drift, and we want to find the cause for that. (Note the subtle difference that we are now looking for the causal parents of the drift of $Y$, and not all causal parents of $Y$). The underlying, unknown, cause is that a certain component of the machine is degrading over time. Within our setting this corresponds, for example, to the parameter $\beta^{e}_{1}$ changing over time. We chose the environments as data from different time intervals, such that the local linearity condition is approximately true. In that case Assumption 1 is (approximately) fulfilled if time affects the drift only through the degradation of $\beta^{e}_{1}$.
3.3 Model strengths and limitations
As our prescribed default goal is to find all causal parents, a problem one may face is that some causal parents and the corresponding structural parameters are not heterogeneous enough across environments. Consider again the disease example and assume the diet was indeed a cause for the disease. If we collect data only from countries with the same diet and its structural parameter (i.e. the way the diet affects the disease) is always the same, LoLICaP will not find diet as a causal parent. From the perspective of environments as interventions it is clear that we are not intervening on the diet, and therefore can generally not draw causal conclusions about this covariate.
Another potential problem is that there is a causal parent which is heterogeneous in the environments, but which we do not observe. This corresponds to the violation of the assumption that the unobserved noise has a fixed distribution, i.e., $\varepsilon^{e}_{i}\sim F^{*}$ for all $e\in[E]$ and all $i\in[n^{e}].$ From the causal perspective this can be seen as the problem that the intervention is a confounder of the target.
4 On the identifiability of causal parents
Recall that, ultimately, we want to build a method based on observational data that is able to identify a set $\tilde{S}\subseteq[D]$ of causal parents. This set desirably approximates $S^{*}$ as well as possible, and ideally $\tilde{S}=S^{*}$. This section takes steps leading to a (meta)-methodology that ensures that $\tilde{S}\subseteq S^{*}$ with high probability, so that we guarantee at least control over false discoveries. This approach is rooted on a testing perspective, and aims to identify sets of plausible causal parents, which are also often called invariant sets (of covariates) in the relevant literature. Roughly speaking, a subset $S\subseteq[D]$ is plausible if it allows for a data generation model as described by Assumption 1, when $S$ takes the place of $S^{*}$. In that sense the name invariant set refers to the property that the covariates with indices in $S$ can model the target $Y$ with an additive noise $\varepsilon^{e}$ that is invariant in $e$. Importantly, the true parents $S^{*}$ are an invariant set by Assumption 1, so finding a minimal invariant set of some sorts allows us to control the false positive probability.
4.1 Control of false positives
In this subsection we define a null hypothesis for each subset $S\subseteq[D]$, which will help us decide if $S$ is a set of plausible causal parents:
$\tilde{H}_{0,S}:\begin{cases}&\text{there exists a distribution $F$ and $\gamma^{e}\in\mathbb{R}^{|S|}$, s.t. $\forall e\in[E]$:}\\ &\text{$Y^{e}=X^{e}_{S}\gamma^{e}+r^{e}$ and for all $i,j\in[n^{e}],i\neq j$:}\\ &\text{$r^{e}_{i}\sim F,r^{e}_{i}\perp\!\!\!\perp r^{e}_{j},r^{e}_{i}\perp\!\!\!\perp X^{e}_{S,i}$}\ .\end{cases}$
We note that $\tilde{H}_{0,S}$ corresponds essentially to Assumption 1 in relation to set $S$. With that we know in particular that, under Assumption 1, $\tilde{H}_{0,S^{*}}$ is true. While $\tilde{H}_{0,S}$ has an exact correspondence to Assumption 1, it is, however, not clear how to build a practical test based on it, that is also powerful against meaningful alternatives. For that reason we move towards a weaker but more practical formulation. Let
$\tilde{\beta}^{e}_{S}=\mathbf{E}[(X^{e}_{S})^{t}X^{e}_{S}]^{\dagger}\mathbf{E}[(X^{e}_{S})^{t}Y^{e}],$
where $A^{\dagger}$ denotes the generalized Moore-Penrose inverse of a matrix $A$ (Penrose, 1955), which always exists and is unique, and where we set $\tilde{\beta}^{e}_{S}=0$ for $S=\emptyset$. With this we can formulate a simpler null hypothesis as
$H_{0,S}:\begin{cases}&\text{there exists a distribution $F$ such that for all $e\in[E]$:}\\ &\text{$Y^{e}=X^{e}_{S}\tilde{\beta}^{e}_{S}+r^{e}$ and for all $i\in[n^{e}]:r^{e}_{i}\sim F$\ .}\end{cases}$
The following lemma, proven in Appendix A, establishes the relation between $H_{0,S}$ and $\tilde{H}_{0,S}$:
Lemma 1.
If $\tilde{H}_{0,S}$ is true then so is $H_{0,S}$.
Remark 1.
The definition of the null hypothesis ${H}_{0,S}$ already foreshadows the relevance of the different environments. Consider the case that we only have one environment $e_{1}$ and the rows of $X^{e_{1}}$ are i.i.d (independently and identically distributed) samples from the same distribution. In that case even $S=\emptyset$ fulfills ${H}_{0,S}$ for $\gamma^{e_{1}}$ being the null-vector and $r^{e}=\beta^{e_{1}}X^{e_{1}}+\varepsilon^{e_{1}}$. This is in accordance with our intuition of thinking about the environments as interventions. If we only have a single intervention, but no control, we cannot draw causal conclusions about the effect of that intervention.
Based on the hypothesis $H_{0,S}$ we are going to consider the following meta-procedure: Recall that $(\boldsymbol{X},\boldsymbol{Y})=\{(X^{e},Y^{e})\}_{e\in[E]}$ are our observations, and let $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})\in\{0,1\}$ be a test function, such that $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=1$ indicates we reject ${H}_{0,S}$. We know that $\tilde{H}_{0,S^{*}}$ holds by Assumption 1, and thus Lemma 1 implies $H_{0,S^{*}}$ also holds, and we expect that with high probability $\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=0$. With this in mind we view all $S$ for which $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0$ as a plausible set, and naturally define the estimator $\tilde{S}$ of $S^{*}$ as
$\tilde{S}\vcentcolon=\bigcap_{S:\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0}S.$
Based on the above considerations we have the following result:
Proposition 1.
Let $\alpha\in(0,1)$. Consider a class of test functions $\phi_{S}$ for all $S\subseteq[D]$. Assume this class of tests satisfies $P[\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=1\mid{H}_{0,S^{*}}\text{\ \ holds}]\leq\alpha$. Then under Assumption 1 we have $\tilde{S}\subseteq S^{*}$ with probability of at least $1-\alpha$.
This result, which is essentially Theorem 1 from Peters et al. (2016), is a simple consequence of the fact that $P[\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=1\mid{H}_{0,S^{*}}\text{\ \ holds}]\leq\alpha$. Note that for this result it is only important to guarantees the good behavior of the testing procedure for the true set $S^{*}$.
4.2 Control of false negatives
The previous section shows that we can guarantee that $\tilde{S}\subseteq S^{*}$ with arbitrarily high probability. However, such a guarantee can be trivially obtained for the choice $\tilde{S}=\emptyset$. A natural question to ask is when is it possible to have a class of tests that ensure that $\tilde{S}=S^{*}$ with high probability. Naturally the answer to this question is more intricate, and depends crucially on how much information is present on data. For the following result we focus on scenarios where one has an arbitrarily large amount of data in each environment, and identify conditions under which the methodology above might be able to identify $S^{*}$ as the only possible set of plausible causal parents.
Specifically, suppose one has access to $\tilde{\beta}^{e}_{S}$ for any $e\in[E]$ and $S\subseteq[D]$. In practical terms this would essentially require to observe the randomly generated data in all environments an infinite number of times. Given access to $\tilde{\beta}^{e}_{S}$, which requirements on the heterogeneity of the environments are needed to ensure one can pinpoint $S^{*}$ as the only set of plausible causal parents? Within a fairly general class of structural equation models for the covariates $X^{e}$, heterogeneity is almost always guaranteed through the changing parameters $\beta^{e}$, and only in rather pathological combinations of parameters problems will arise. This is made precise in Theorem 1. The class of structural equation models for which we can show identifiability is given through the following assumption:
Assumption 2.
To simplify notation, we introduce the index $y\vcentcolon=D+1$ and define $X^{e}_{y}\vcentcolon=Y^{e}$. We assume that $E\geq 2$ and for all $e\in E$ it holds that $n^{e}>0$. Furthermore for at least one $e\in E$ there exists an acyclic graph $G=([D]\cup{y},\mathcal{E})$ such that for $d\in[D]$ the structural equation of $X^{e}_{d}$ has the form
$X^{e}_{d}\vcentcolon=f_{d}(X^{e}_{\textbf{PA}(d)})+\delta^{e}_{d},$
where $f_{d}$ is a polynomial of finite degree in $X^{e}_{y}=Y^{e}$ if $y\in\textbf{PA}(d)$, but otherwise arbitrary. Here $\delta^{e}_{d}=(\delta^{e}_{d,1},\cdots,\delta^{e}_{d,n^{e}})$ is a random noise vector such that
$\delta^{e}_{d}\sim\mathcal{D}^{e}_{d}$
$\forall:i\in[n^{e}]\quad\delta^{e}_{d,i}\perp\!\!\!\perp X^{e}_{u,i}\quad\text{\ for all \ }u\in\textbf{NDE}(d),$ (4)
where $\mathcal{D}^{e}_{d}$ is a distribution such that $\Delta^{e}_{d,i}\vcentcolon=\mathbf{V}(\delta^{e}_{d,i})>0$ for all $i\in[n^{e}]$. We define $\Delta^{e}\in(0,\infty)^{D\times n^{e}}$ to be the matrix with the $(d,i)$-th entry given by the variance $\Delta^{e}_{d,i}$. Finally, we assume that for all $e\in[E],d\in[D]$ and $i\in[n^{e}]$ the covariate $X^{e}_{d,i}$ has finite variance.
Informally speaking the noise terms $\delta_{d}^{e}$ ensure that each covariate introduces unique information and prevent that the causal parents $X^{e}_{S^{*}}$ lie in the column space of any other subset of variables $X^{e}_{S}$ with $S^{*}\not\subseteq S$. The need for the explicit variances $\Delta^{e}_{d,i}$ is due to our proof technique, and we contextualize this assumption further after we state the theorem. The polynomial dependence on $Y^{e}$ simplifies our proof, could, however, be replaced by other regularity assumptions on $f_{d}$. With this in hand we are ready to state our main identifiability result:
Theorem 1.
Let $S\subset[D]$ such that $S^{*}\not\subseteq S$ and take any two environments $e,v\in[E]$ with $e\neq v$, such that environment $e$ fulfills the data generation mechanism from Assumption 2 with variances $\Delta^{e}$. Suppose Assumption 1 holds with parameters $\beta^{e},\beta^{v}$. Then there is a set $M_{0}\subset\mathbb{R}^{|S^{*}|\times|S^{*}|}\times(0,\infty)^{D\times n^{e}}$ with Lebesgue measure zero, such that if
$(\beta^{v},\beta^{e},\Delta^{e})\in\mathbb{R}^{|S^{*}|\times|S^{*}|}\times(0,\infty)^{D\times n^{e}}\setminus M_{0}$
it is guaranteed that $H_{0,S}$ is false.
Proof.
The general proof strategy is the following: From the two distinct environments $e,v\in[E]$ we pick a sample $i\in[n^{e}]$ and $j\in[n^{v}]$. If the hypothesis $H_{0,S}$ would be true we may conclude that the population residuals $r^{e}_{i}$ and $r^{v}_{j}$ have the same distribution, and therefore the same variance. We then show, however, that there exists a $u\in S^{*}$ and $u_{1}\in[D]$ such that $\mathbf{V}[r^{e}_{i}]$ is a proper rational function (so the ratio of two polynomials) of finite degree with respect to $\beta^{e}_{u}$ for almost all choices of $\Delta^{e}_{u_{1},i}$. Fixing all entries of $(\beta^{e},\beta^{v},\Delta^{e})$ except $\beta^{e}_{u}$ and $\Delta^{e}_{u_{1},i}$ at arbitrary values, we conclude that the equation $\mathbf{V}[r^{e}_{i}]=\mathbf{V}[r^{v}_{j}]$ can be solved for at most finitely many values of $\beta^{e}_{u}$ and $\Delta^{e}_{u_{1},i}$. This means that $M_{0}$, the solution space of the equation $\mathbf{V}[r^{e}_{i}]=\mathbf{V}[r^{v}_{j}]$ with respect to $(\beta^{e},\beta^{v},\Delta^{e})$, has Lebesgue measure zero. This finally implies that the hypothesis $H_{0,S}$ is false for all parameter choices outside of $M_{0}$.
To start the proof, we note that $\mathbf{V}[r^{e}_{i}]$ is always a rational function of polynomials of finite degree with respect to the entries of $\beta^{e}$. On the one hand this follows from Assumption 2, ensuring that every covariate $d$ with $y\in\textbf{AN}(d)$ is a polynomial in $Y^{e}$ and thus a polynomial in $\beta^{e}$. On the other hand we know from Constales (1998) that the Moore-Penrose inverse has a closed form solution consisting only of elementary operations. The rest of the proof is concerned with showing that there is at least one $u\in S^{*}$ such that $\mathbf{V}[r^{e}_{i}]$ is a proper polynomial with respect to $\beta^{e}_{u}$, meaning that the leading term in $\beta^{e}_{u}$ does not vanish. To simplify notation, we define
$P(S,S^{*})\vcentcolon=\mathbf{E}[(X^{e}_{S})^{t}X^{e}_{S})]^{\dagger}\mathbf{E}[(X^{e}_{S})^{t}X^{e}_{S^{*}}]\in\mathbb{R}^{|S|\times|S^{*}|}$
and
$P(\varepsilon^{e})\vcentcolon=\mathbf{E}[(X^{e}_{S})^{t}X^{e}_{S})]^{\dagger}\mathbf{E}[(X^{e}_{S})^{t}\varepsilon^{e}]\in\mathbb{R}^{|S|}$
and note that $\tilde{\beta}^{e}_{S}=P(S,S^{*})\beta^{e}_{S^{*}}+P(\varepsilon^{e})$. We split the proof into two cases.
The first case assumes that no variable in $S$ is a descendant of $Y$. In that case we pick any $u\in S^{*}\setminus S$ and split the variance of the residual as follows:
$\mathbf{V}[r^{e}_{i}]$
$=\mathbf{V}[X^{e}_{S^{*},i}\beta^{e}_{S^{*}}+\varepsilon^{e}_{i}-X^{e}_{S,i}\tilde{\beta}^{e}_{S}]=\mathbf{V}[(X^{e}_{S^{*},i}-X^{e}_{S,i}P(S,S^{*}))\beta^{e}_{S^{*}}+\varepsilon^{e}_{i}-X^{e}_{S,i}P(\varepsilon^{e})]$
$=\mathbf{V}\Biggl{[}(X^{e}_{u,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,u})\beta^{e}_{u}$
$\quad+\sum_{d\in S^{*}\setminus\{u\}}(X^{e}_{d,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,d})\beta^{e}_{d}+\varepsilon^{e}_{i}-X^{e}_{S,i}P(\varepsilon^{e})\Biggr{]}$
$=(\beta^{e}_{u})^{2}\mathbf{V}[(X^{e}_{u,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,u})]$ (5)
$\quad+\mathbf{V}\left[\sum_{d\in S^{*}\setminus\{u\}}(X^{e}_{d,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,d})\beta^{e}_{d}+\varepsilon^{e}_{i}-X^{e}_{S,i}P(\varepsilon^{e})\right]$ (6)
$\quad+2\beta^{e}_{u}\mathbf{C}\left[(X^{e}_{u,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,u})\ ,\sum_{d\in S^{*}\setminus\{u\}}(X^{e}_{d,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,d})\beta^{e}_{d}+\varepsilon^{e}_{i}-X^{e}_{S,i}P(\varepsilon^{e})\right]\ .$ (7)
We now argue that the variance in line (5) does not vanish (Claim 1 further below), and that the covariance in line (7) is only linear in $\beta^{e}_{u}$ (Claim 2 further below). With that, and noting that the variance term in line (6) is always positive, we know that there are constants $a>0$ and $b\in\mathbb{R}$ (which depend on the distributions of the covariates) such that
$\mathbf{V}[r^{e}_{i}]\geq(\beta^{e}_{u})^{2}a+\beta^{e}_{u}b.$ (8)
With that $\mathbf{V}[r^{e}_{i}]$ is a proper polynomial in $\beta^{e}_{u}$, as $\mathbf{V}[r^{e}_{i}]\to\infty$ for $\beta^{e}_{u}\to\infty$.
Claim 1.
The variance in line (5) does not vanish. Proof: Let $\bar{S}\subset S$ be the set of all indices $d\in S$ with $P(S,S^{*})_{d,u}\neq 0$. All indices outside $\bar{S}\cup\{u\}$ are irrelevant as the corresponding covariate vanishes within the variance term (5). Now let $u_{0}$ be a sink node in $\bar{S}\cup\{u\}$, which implies by Assumption 2 that $\Delta^{e}_{u_{0},i}$ is independent of all other covariates with index in $\bar{S}\cup\{u\}$. First, let $u_{0}\in\bar{S}$, then we may split the variance from line (5) as
$\mathbf{V}[(X^{e}_{u,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,u})]$
$=\mathbf{V}[X^{e}_{u,i}-X^{e}_{S,i}P(S,S^{*})_{\cdot,u}+\delta^{e}_{u_{0},i}P(S,S^{*})_{u_{0},u}]+\mathbf{V}[\delta^{e}_{u_{0},i}]P(S,S^{*})^{2}_{u_{0},u}.$
This splitting is allowed as by design $\delta^{e}_{u_{0},i}$ is independent of all covariates $X^{e}_{d,i}$ within the variance with $d\neq u_{0}$, and the dependence on $X^{e}_{u_{0},i}$ is canceled by the term $+\delta^{e}_{u_{0},i}P(S,S^{*})_{u_{0},u}$. From the definition above we know that $P(S,S^{*})_{u_{0},u}\neq 0$ and by Assumption 2 we have $\mathbf{V}[\delta^{e}_{u_{0},i}]>0$, which implies Claim 1. The case that $u_{0}=u$ follows analogously by splitting $\delta^{e}_{u,i}$ out of the variance.
Claim 2.
The covariance in line (7) is only linear in $\beta^{e}_{u}$. Proof: This is a direct consequence from the fact that no covariate is a descendant of $Y^{e}$, which also implies that no covariate has a dependence on $\beta^{e}_{u}$. The covariance does then not depend on $\beta^{e}_{u}$, and we only have a linear dependency in line (7) from the leading coefficient.
The second case assumes there exists at least one $d\in S$ with $d\in\textbf{DE}(y)$, where by acyclicity of $G$ we know that $d\not\in S^{*}$. Let $u\in S^{*}\setminus S$ and we fix all entries of $(\beta^{e},\beta^{v})$, except $\beta^{e}_{u}$. We now show that there is an index $u_{1}\in[D]$ such that for almost all values of $\Delta^{e}_{u_{1},i}$ the variance term $\mathbf{V}[r^{e}_{i}]$ is different for $\beta^{e}_{u}=0$ and $\beta^{e}_{u}\to\infty$. This implies that $\mathbf{V}[r^{e}_{i}]$ is for almost all values of $\Delta^{e}_{u_{1},i}$ a ratio of proper polynomial in $\beta^{e}_{u}$ and the argumentation follows as in the first case. To make this argumentation formal we use the following two claims:
Claim 3.
For $\beta^{e}_{u}=0$ and any $u_{0}\in\textbf{DE}(y)$ the term $\mathbf{V}[r^{e}_{i}]$ is finite for $\Delta^{e}_{u_{0},i}\to\infty$. Proof: We just have to note that $u_{0}\not\in S^{*}$, which implies that $\mathbf{V}[Y^{e}_{i}]$ remains finite for $\Delta^{e}_{u_{0},i}\to\infty$. The claim follows by noting that $\mathbf{V}[r^{e}_{i}]\leq\mathbf{V}[Y^{e}_{i}]+\mathbf{E}[Y^{e}_{i}]^{2}$.
Claim 4.
For all $u_{0}\in\textbf{DE}(y)$ set $\Delta^{e}_{u_{0},i}=q$, then for $q\to\infty$ the term $\mathbf{V}[r^{e}_{i}]$ diverges for $\beta^{e}_{u}\to\infty$. Proof: We may assume that there exists at least one $u_{0}\in\textbf{DE}(y)$ such that $\tilde{\beta}^{e}_{u_{0}}\neq 0$ for $q\to\infty$, otherwise we are back to the first case as we then may remove all descendants of $y$ from our set $S$. In this first case we have shown Claim 4 already after Inequality (8). Without loss of generality let $u_{0}\in\textbf{DE}(y)$ be a sink node, which implies that $\delta_{u_{0}}$ is independent of all other variables by (4). Then we may split $\mathbf{V}[r^{e}_{i}]$ as
$\mathbf{V}[r^{e}_{i}]$ $=\mathbf{V}[Y-X^{e}_{S}\tilde{\beta}^{e}_{S}+\delta^{e}_{u_{0},i}\tilde{\beta}^{e}_{u_{0}}-\delta^{e}_{u_{0},i}\tilde{\beta}^{e}_{u_{0}}]$
$=\mathbf{V}[Y-X^{e}_{S}\tilde{\beta}^{e}_{S}+\delta^{e}_{u_{0},i}\tilde{\beta}^{e}_{u_{0}}]+\mathbf{V}[\delta^{e}_{u_{0},i}](\tilde{\beta}^{e}_{u_{0}})^{2}.$
Since by assumption $\tilde{\beta}^{e}_{u_{0}}\neq 0$ we observe that $\mathbf{V}[r^{e}_{i}]\to\infty$ for $\mathbf{V}[\delta^{e}_{u_{0},i}]=\Delta^{e}_{u_{0},i}\to\infty$.
The above Claim 3 and Claim 4 together imply that $\mathbf{V}[r^{e}_{i}]$ obtain different values for $\beta^{e}_{u}=0$ and $\beta^{e}_{u}\to\infty$ in the regime that $\Delta^{e}_{u_{0},i}\to\infty$ for all $u_{0}\in\textbf{DE}(y)$. This implies that in this regime the term $\mathbf{V}[r^{e}_{i}]$ is a proper rational function in $\beta^{e}_{u}$. It could, however, still happen that for specific choices of $\Delta^{e}_{\cdot,i}\in(0,\infty)^{D}$ the rational function dependence on $\beta^{e}_{u}$ cancels within $\mathbf{V}[r^{e}_{i}]$. More precisely, let $c$ be the leading coefficient of the polynomial term in $\beta^{e}_{u}$, then it is still possible that $c=0$ for specific choices of $\Delta^{e}_{\cdot,i}$ as $c$ generally depends on those terms. However, let $u_{1}\in[D]$ be any index such that $c$ depends on $\Delta^{e}_{u_{1},i}$. Fixing all entries in $\Delta^{e}_{\cdot,i}$ except $\Delta^{e}_{u_{1},i}$ we know that $c=0$ for at most finitely many choices of $\Delta^{e}_{u_{1},i}$ since $\mathbf{V}[r^{e}_{i}]$ is also a rational function in $\Delta^{e}_{u_{1},i}$.
We thus have shown that there exists a $u\in S^{*}\setminus S$ such that $\mathbf{V}[r^{e}_{i}]$ is a proper rational function in $\beta^{e}_{u}$ for almost all values of $\Delta^{e}_{\cdot,i}$. With that we know that for almost all values of $\Delta^{e}_{\cdot,i}$ the equation $\mathbf{V}[r^{e}_{i}]=\mathbf{V}[r^{v}_{i}]$ can be solved for at most finitely many choices of $\beta^{e}_{u}$. With that, the solution space $M_{0}$ of the equation $\mathbf{V}[r^{e}_{i}]=\mathbf{V}[r^{v}_{i}]$ with respect to the parameters $(\beta^{e},\beta^{v},\Delta^{e})$ has a Lebesgue measure of zero. As the equality of $\mathbf{V}[r^{e}_{i}]$ and $\mathbf{V}[r^{v}_{i}]$ is a necessary condition for $H_{0,S}$ to be true, we know that $H_{0,S}$ is false for all parameter values outside of $M_{0}$.
It is important to emphasize that the inclusion of $\Delta^{e}_{d,i}$ as additional parameters is an operational way to ensure each covariate introduces unique information. Similar results can be attained by making different sets of assumptions achieving the same goal. Also note that we do not require $\Delta^{e}_{d,i}$ to be heterogeneous in the environments. For instance, we do not exclude a scenario where $\Delta^{e}_{d,i}$ is the same for all $e\in[E]$, or have any other type of relationship across environments. The same does not hold for $\beta^{e}$ and $\beta^{v}$ as they both appear in $M_{0}$.
The role of local linearity.
At this point we want to make an interesting observation following the previous analysis. In the work of Peters et al. (2016) the heterogeneity across environments is always introduced through (accidental) interventions on the covariate distributions, and this in turn allows them to reject a hypothesis similar to $H_{0,S}$ if $S^{*}\not\subseteq S$. In contrast to that our Theorem 1 states that even if the covariates $X^{e}$ have all the same distributions for different $e\in[E]$, the hypothesis $H_{0,S}$ is still not true for almost all structural parameters $\beta^{e}$ if $S^{*}\not\subseteq S$. In that sense, the assumption of local linearity not only makes our model more flexible, but also introduces a more general concept of heterogeneity.
Recovering the full causal graph.
Another interesting observation following Theorem 1 is that within the proposed setting, and under further assumptions, one can recover the full causal graph, and not only the causal parents of a chosen target. In particular we can switch the role of $Y^{e}$ with any other covariate $X^{e}_{d}$, if the environment index $e$ is not a direct causal parent of $X^{e}_{d}$, in which case Assumption 1 remains true. We then still have identifiability of the causal parents of $X^{e}_{d}$ under the assumptions of Theorem 1. In Section 6.3 we present experimental evidence for this use case.
5 Proposed Approach and Finite Sample Results
The statement of Theorem 2 relies on the values $\tilde{\beta}^{e}_{S}$, which are not observable. That being said, the proposed methodological idea can be made practical by replacing $\tilde{\beta}^{e}_{S}$ by suitable estimates based on the observed data. In this section we make this concrete, and also provide a finite-sample characterization of the proposed procedure in a simplified scenario.
The first step is to identify a suitable surrogate for $\tilde{\beta}^{e}_{S}$ based on the observations $(\boldsymbol{X},\boldsymbol{Y})$. A natural candidate is simply the solution of a generalized least-squares problem:
$\hat{\beta}^{e}_{S}=((X^{e}_{S})^{T}X^{e}_{S})^{\dagger}(X^{e}_{S})^{T}Y^{e},$
where we set $\hat{\beta}^{e}_{S}$ as the null vector when $S=\emptyset$. Based on the estimates $\hat{\beta}^{e}_{S}$ we also define the corresponding residuals
$r^{e}_{S}={Y}^{e}-{X}^{e}_{S}\hat{\beta}^{e}_{S}.$
Recall that the methodological idea is to define the estimator $\tilde{S}$ of $S^{*}$ by setting
$\tilde{S}\vcentcolon=\bigcap_{S:\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0}S,$
where $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})$ is a test for the hypothesis $H_{0,S}$ based on data $(\boldsymbol{X},\boldsymbol{Y})$. Therefore, the second step is to construct a test $\phi_{S}$. To facilitate this, we make the following Gaussian noise assumption.
Assumption 3.
The random noise variables ${\varepsilon}^{e}_{i}$ are sampled from a Gaussian distribution with zero mean and unknown variance $\sigma^{2}_{Y}$. Furthermore, we assume the following independencies for all $e,v\in[E]$ with $e\neq v$
$\varepsilon^{e}_{i}\perp\!\!\!\perp\varepsilon^{v}_{j}\text{ \ \ for all \ \ }i\in[n^{e}],j\in[n^{v}]$
$\varepsilon^{e}_{i}\perp\!\!\!\perp\varepsilon^{e}_{j}\text{ \ \ for all \ \ }i,j\in[n^{e}],i\neq j.$
While this is a strong distributional assumption on the observation noise, the proposed methodology will still be approximately sound when the number of observations in each environment is large (in view of the central limit theorem). However, this assumption facilitates the analysis by not introducing unnecessary clutter. In the experimental section we further examine the robustness of the methodology under violation of the normality assumption.
With all the ingredients in hand, we propose a test based on the following statistic:
$T_{S}(\boldsymbol{X},\boldsymbol{Y})\vcentcolon=\left\{\begin{array}[]{ll}\min_{e\in[E]}\|r^{e}_{S}\|^{2}_{2}/\max_{e\in[E]}\|r^{e}_{S}\|^{2}_{2}&\text{ if }\quad\exists e\in[E]:\|r^{e}_{S}\|_{2}>0\\ \infty&\text{ otherwise}\end{array}\right..$
Under a Gaussian noise assumption, we have the useful property that ${\sigma^{2}_{Y}}\|r^{e}_{S^{*}}\|^{2}_{2}$ are all chi-squared distributed (note we are considering $S^{*}$), and the number of degrees of freedom depends only on the properties of the Gram matrix. Importantly, the scaling $\sigma^{2}_{Y}$ is the same for all $e\in[E]$, which implies that the distribution of $T_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})$ is not a function of $\sigma^{2}_{Y}$. Therefore, we can calibrate a test based on $T_{S}(\boldsymbol{X},\boldsymbol{Y})$ using only on observed quantities. This test statistic is motivated by the problems of sparse testing (Ingster, 1997; Donoho and Jin, 2004; Stoepker et al., 2021), and it targets scenarios where we expect evidence for rejection of the null hypothesis is present on a relatively small number of environments.
To calibrate a test based on this statistic we proceed as follows: For $e\in[E]$ define $Z^{e}_{S}$ as mutually independent chi-squared random variables, respectively with $n^{e}-\text{rank}((X_{S}^{e})^{T}X_{S}^{e})$ degrees of freedom (zero degrees of freedom correspond to $Z^{e}_{S}=0$). Given $(\boldsymbol{X},\boldsymbol{Y})$ these variables are also independent of all the other quantities. Finally, we define the test $\phi_{S}$ as
$\phi_{S}(\boldsymbol{X},\boldsymbol{Y},\alpha)=\left\{\begin{array}[]{ll}1&\text{ if }P\left(\left.T_{S}(\boldsymbol{X},\boldsymbol{Y})>\frac{\min_{e\in[E]}Z^{e}_{S}}{\max_{e\in[E]}Z^{e}_{S}}\right|\boldsymbol{X},\boldsymbol{Y}\right)\leq\alpha\\ 0&\text{ otherwise}\end{array}\right..$ (9)
Note that under the Gaussian assumption this test is well calibrated. Strictly speaking to run the test we need to know the distribution of ${\min_{e\in[E]}Z^{e}_{S}}/{\max_{e\in[E]}Z^{e}_{S}}$. This is not easy to characterize analytically, but we can very easily generate samples from it, so calibration by Monte-Carlo simulation is extremely easy and convenient. The overall procedure, called LoLICaP, is described in Algorithm 3. Theorem 1 then suggest that for any set $S$ with $S^{*}\not\subseteq S$ the distribution of $\|r^{e}_{S}\|^{2}_{2}$ will not be a scaled chi-squared, and so the above statistic will be sensitive to such a change. Whether or not we can actually detect this difference in behavior depends on many factors such as the sample size, strength of the noise, how heterogeneous the environments are, and so on.
{Algorithm}
Input: $(\boldsymbol{X},\boldsymbol{Y}),\alpha$. In order: observations, confidence level Output: $\tilde{S}$, the estimated causal parents For all $S\subseteq[D]$ Set $\hat{\beta}^{e}_{S}\vcentcolon=((X^{e}_{S})^{T}X^{e}_{S})^{\dagger}(X^{e}_{S})^{T}Y^{e}$ for all $e\in[E]$ Set $r^{e}_{S}\vcentcolon=Y^{e}-X_{S}^{e}\hat{\beta}^{e}_{S}$ for all $e\in[E]$ Set $T_{S}\vcentcolon=\min_{e\in[E]}\|r^{e}_{S}\|^{2}_{2}/\max_{e\in[E]}\|r^{e}_{S}\|^{2}_{2}$ Set $\tilde{S}\vcentcolon=\bigcap\limits_{S:\phi_{S}(\boldsymbol{X},\boldsymbol{Y},\alpha)=0}S$ Our proposed method LoLICaP.
5.1 Controlling False Positives
The correct coverage properties are a direct consequence of the work we did in the previous subsection:
Proposition 2.
Consider Assumptions 1 and 3 , and let $\alpha\in(0,1)$. Then
$\mathbf{P}(\tilde{S}\not\subseteq S^{*})\leq\alpha,$
for $\tilde{S}$ being the output of Algorithm 3.
Proof.
If for all $e\in[E]$ the matrix $(X^{e}_{S^{*}})^{T}X^{e}_{S^{*}}$ is not invertible we are done, as in that case $\tilde{S}=\emptyset$. If this matrix is invertible, then we know by Assumptions 1 and 3 that ${r^{e}_{S^{*}}\sim\sigma^{2}_{Y}\chi^{2}(n^{e}-\text{rank}((X_{S^{*}}^{e})^{T}X_{S^{*}}^{e}))}$. With that $T_{S^{*}}=\min_{e\in[E]}\|r^{e}_{S^{*}}\|_{2}^{2}/\max_{e\in[E]}\|r^{e}_{S^{*}}\|_{2}^{2}$ and ${\min_{e\in[E]}Z^{e}_{S}}/{\max_{e\in[E]}Z^{e}_{S}}$ follow the same distribution by definition of $Z^{e}_{S}$, which implies that
[…]
By definition of $\phi_{S^{*}}$ this means that $\mathbf{P}(\phi_{S^{*}}=1)\leq\alpha$, which by definition of $\tilde{S}$ finally implies that $\mathbf{P}(\tilde{S}\not\subseteq S)\leq\alpha.$
The probability of including a false positive parent is relatively easy to understand as this does not depend on anything other than Assumptions 1 and 3. Controlling false negative discoveries, so controlling the probability that $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0$ for $S\subseteq[D]$ with $S^{*}\not\subseteq S$, becomes much more complex. To shed some light into this the next section considers a simplified setting yielding some useful insights.
5.2 False Negative Rate for Independent Normal Covariates
While the previous section ensures that our test is well calibrated, we would also like to understand the probability of missing a true causal parent, i.e., understand the false negative rate. This is much more complex than controlling the false positives: On the one hand the probability for missing a causal parent depends on all kind of quantities as the sample size, the number of environments, the strength of the heterogeneity and more. On the other hand, we do generally not even know the distribution of the sum of squared residuals (SSR) $\|r^{e}_{S}\|^{2}_{2}$ when $S\neq S^{*}$, which makes it difficult to understand the behavior of the test statistics $T_{S}$. To gain control over those quantities we present our results for the finite sample case in a simplified setting, where all covariates are generated as independent normal random variables. In the experimental section we then test how the performance is affected if this assumption is violated in different ways. For now, we make the following specific data generating assumption.
Assumption 4.
Let the number of environments $E$ be even, and let $[E_{1}],[E_{2}]$ denote two index sets for two types of environments such that $[E]=[E_{1}]\dot{\cup}[E_{2}]$ with $|[E_{1}]|=|[E_{2}]|=\frac{E}{2}$. In each individual environment we observe $n>D$ observations (i.e., $\forall e\in[E]\ n^{e}=n$). For all $v\in[E_{1}]$ and $d\in[D]$ the $d-th$ covariate ${X}^{v}_{d}\in\mathbb{R}^{n}$ is an $n$-sample from $N(\mu^{v}_{d},\sigma^{v}_{d})$, sampled independently from each other and the other covariates of the environment $v$. Similarly, for all $w\in[E_{2}]$ we sample ${X}^{w}_{d}$ from $N(\mu^{w}_{d},\sigma^{w}_{d})$ with the same independence assumptions. We assume that there exists $\beta^{1},\beta^{2}\in\mathbb{R}^{D}$ such that $\beta^{v}=\beta^{1}$ and $\beta^{w}=\beta^{2}$ for all $v\in[E_{1}],w\in[E_{2}]$.
Remark 2.
While in the setting above inverse matrices $({X^{w}_{S}}^{T}X^{w}_{S})^{-1}$ and $({X^{v}_{S}}^{T}X^{v}_{S})^{-1}$ exist for $v\in[E_{1}]$ and $w\in[E_{2}]$ and any $S\subseteq[D]$ with probability one, we consider for simplicity only the case that they exist.
While the independent normal assumption is a strong restriction on the covariates, it is already a meaningful setting in the sense that without further assumptions one cannot distinguish cause and effect in this case, as seen for example in Peters and Bühlmann (2014). The additional assumption we make are the heterogeneous environments. Missing a causal parent so that $S^{*}\subsetneq\tilde{S}$ happens if and only if there exists a set $S\subseteq[D]$ with $S^{*}\not\subseteq S$ that we accept as a plausible set, so that $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0$. To understand the distribution of $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})$ we need a notion of how much environment heterogeneity is introduced by the covariates in $U\vcentcolon=S^{*}-S$, as this is the driver of LoLICaP to identify causal parents. In this simplified setting the heterogeneity is carried by the following ratio:
$I_{S}=\min\left\{\frac{\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}}{\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}},\frac{\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}}{\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}}\right\},$
where $(\sigma_{Y})^{2}$ is the variance of the target noise $\varepsilon^{e}$. Note that $0<I_{S}\leq 1$ and small values of $I_{S}$ indicate a higher environment heterogeneity. Note in particular that $I_{S}=1$ if for all $u\in U$ we have $(\beta^{1}_{u})^{2}=(\beta^{2}_{u})^{2}$ and $(\sigma^{v}_{u})^{2}=(\sigma^{w}_{u})^{2}$. The following result presents bounds on the false negative probability in terms of the sample size $n$ and the heterogeneity parameter $I_{S}$. We already disclaim that the result treats the effect of the number of environments $E$ crudely, and due to our proof technique it is actually vacuous for the case that $E\to\infty$. We instead chose to analyze the setting $E\to\infty$ in isolation, and a corresponding result is presented afterwards.
Theorem 2.
For $S\subset[D]$, with $S^{*}\not\subseteq S$ define $I_{S}$ as above and set $k\vcentcolon=n-|S|$. If Assumptions 1,3, 4 are true and $I_{S}<1$, then for any confidence level $\alpha\geq 0$ it holds that
$\mathbf{P}(\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0)\leq\frac{4E}{\alpha}\left(\left(\frac{1}{(I_{S})^{\frac{1}{4}}}e^{(1-{1}/{{(I_{S})^{\frac{1}{4}}}})}\right)^{\frac{k}{2}}+\left({(I_{S})^{\frac{1}{4}}}e^{(1-{(I_{S})^{\frac{1}{4}}})}\right)^{\frac{k}{2}}\right).$
This means that the probability to accept $S$ falsely as a plausible set drops exponentially fast in $k$, since $(ce^{1-c})<1$ for any $c\neq 1$.
The proof of this result is presented in Appendix A. Before we tackle the case $E\to\infty$ a few points are in order. First, we want to highlight that this result concerns only an individual subset $S\subset[D]$. Since LoLICaP iterates over all such $S$, we perform exponentially many such tests with respect to the dimensionality $D$. One could, of course, derive a version of Theorem 2 that holds for all subsets by means of a union bound, but such a result would be unnecessarily crude: many of those tests are expected to be positively correlated, as different choices of $S$ do overlap. For this reason, we do not attempt to derive such a result.
The next interesting observation following Theorem 2 is a comparison to a similar result from Pfister et al. (2019), see also Figure 4 within that paper. They extend the work of Peters et al. (2016) to sequential data and an analysis of their method yields a Theorem, corresponding to our Theorem 2, with a convergence rate of $O(n^{-\frac{1}{2}})$. The exact reason for this discrepancy is beyond the analysis of this paper, and possibly hard to identify as their methodology and setting differs in many ways from ours after the same starting point of assuming structural invariance. We, however, reiterate the main difference in their setting: They assume that there is a global linear model, and testing if a set of covariates is a plausible set of causal parents includes a test of the global linearity condition. This is combined with a test based on the ratio of the SSR of different environments. As there is a resemblance in that to our own test statistic, it is the more surprising that the resulting convergence rates are quite different.
After those two observations we still owe the reader a result elucidating the case $E\to\infty$:
Theorem 3.
Let $S\subseteq[D]$, with $S^{*}\not\subseteq S$, $I_{S}$ as defined above, and $k=n-|S|$. To emphasize the dependence of the data on $E$ we write now $(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})=\{(X^{e},Y^{e})\}_{e\in E}$. If Assumptions 1,3,4 hold then for any $\alpha\geq 0$ we have that
$\lim\limits_{E\to\infty}\mathbf{P}(\phi_{S}(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})=0)\leq\frac{1}{\alpha}\frac{2(I_{S})^{\frac{k}{2}}}{2(I_{S})^{\frac{k}{2}}+1}.$
If we further assume the collection of random variables $\{X^{e_{1}}_{d_{1},i_{1}},X^{e_{2}}_{d_{2},i_{2}}\}$ for $e_{1}\in[E_{1}],e_{2}\in[E_{2}]$ and $i_{1},i_{2}\in[n]$ to be mutually independent then for any $\alpha<\frac{(I_{S})^{\frac{k}{2}}}{(I_{S})^{\frac{k}{2}}+1}$ it holds that
$\frac{1}{1-\alpha}\left(\frac{(I_{S})^{\frac{k}{2}}}{(I_{S})^{\frac{k}{2}}+1}-\alpha\right)\leq\lim\limits_{E\to\infty}\mathbf{P}(\phi_{S}(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})=0).$
Comparing Theorem 2 and 3 we make the observation that the dependence of the bound on $I_{S}$ and $k$ is very similar, the biggest difference being that Theorem 2 loses a factor of $\frac{1}{4}$ in the exponent of $I_{S}$ compared to Theorem 3. This may, however, very well be due to the proof technique of Theorem 2, in particular the use of Lemma 2. More importantly, Theorem 3 shows that the false acceptance probability does not necessarily converge to 00 for increasing $E$. This indicates a potentially complicated relationship between the number of available environments $E$ and the performance of LoLICaP. To study this relation further, we conduct experiments in the following section.
6 Experimental Results
We now perform a range of experiments to shed further light on the behavior of LoLICaP. The first set of experiments evaluates the performance of the method with respect to the most important quantities such as sample size and number of environments, when all model assumptions hold. Alongside this, we investigate the effects of the violation of some assumptions, in particular the normality assumptions. We observe that in most cases the performance is not crucially affected.
The next range of experiments compares the performance of LoLICaP against a version of LiNGAM that receives data from different environments as input, see also Shimizu (2012). Since there is overlap between the assumptions of both methods, a meaningful comparison is possible. We report that our proposed methodology shows strong performance advantages over LiNGAM, if the environments are sufficiently heterogeneous. By varying the heterogeneity in one of the experiments, we highlight that this is a crucial factor for the performance of our proposed methodology.
Finally, we showcase the use of LoLICaP on a large dataset coming from a dynamical, non-linear, system. This type of data can, for example, be observed in lithography systems, where sensors placed in the machine are sampled at a high frequency. To ensure full control of the evaluation we chose to use artificial data. Nevertheless, this experiment shows that the LoLICaP can find causal parents, also in non-linear (dynamical) systems.
For the implementation of the tests $\phi_{S}$ in all the following experiments we generate $B$ samples from ${\min_{e\in[E]}Z^{e}_{S}}/{\max_{e\in[E]}Z^{e}_{S}}$ to compute the corresponding $p$-values. Unless stated otherwise $B=100$.
6.1 Effects of Non-Normality and Correlated Covariates
We developed our proposed algorithm, as well as the theory in the finite sample case, under normality assumptions of the target noise $\varepsilon^{e}$ and the covariates, i.e., Assumption 4, where covariates are also assumed to be sampled independently from each other. In this section we compare the performance of LoLICaP in this idealized setting against a setting with correlated covariates and non-normal distributions.
While a full description of the data generating processes for the experiments is found in Appendix B, the rough explanation is the following: For the idealized setting we sampled six covariates and the target noise from independent normal random variables, from which two covariates are actual causal parents of the target. We call the data from this process "Normal". Unless stated otherwise we set $E=30$ and in each environment the standard deviations of the covariate distributions as well as the parameters $\beta^{e}$ are independently sampled from a uniform distribution on $[1,5]$. Similarly, we generate data with independent uniformly and Student-t distributed covariates and target noise, thus violating the normality assumptions. Data from those processes are called "Uniform" and "Student-t" respectively. Finally, we generate data from a linear structural equation model with normal noise, which introduces a dependence structure between the covariates. Data from this process is finally called "SEM". We then execute LoLICaP over $300$ independent runs and report how often the method missed a causal parent (false negative rate) and how often the method returned a non-parent (false positive rate). More precisely, let $\tilde{S}_{r}$ be the estimate of the $S^{*}$ in run $r$. The false negative rate is given by $\frac{1}{300}\sum_{r=1}^{300}\mathbf{1}\left\{S^{*}\setminus\tilde{S}_{r}\neq\emptyset\right\}$ and the false positive rate by $\frac{1}{300}\sum_{r=1}^{300}\mathbf{1}\left\{\tilde{S}_{r}\setminus S^{*}\neq\emptyset\right\}$. We also report error bars, which are computed as a $95$ percent Clopper-Pearson confidence interval (Clopper and Pearson, 1934). Unless stated otherwise, all tests of LoLICaP are calibrated at confidence level $\alpha=0.1$.
(a) Normal and SEM data.
(b) Student-t and Uniform data.
Figure 3: We plot the performance of LoLICaP against an increasing number of samples per environment for different distributions of the covariates. Correlation in the data, as introduced by SEM, overall decreases the performance, while for the Student-t distributed data the false positive rate exceeds the set threshold of the confidence level. In all cases, however, the false negative rate behaves similar to the rate observed for the normally distributed data (Normal).
Increasing sample size.
We start by investigating the performance behavior for an increasing sample size per environment. We are interested how the violations to the independence and normality Assumption 4 affects the performance. In Figure 3 we plot the rates of the false positives and false negatives for different sample sizes and the different data generating distributions. While the performance is generally worse in the SEM data, the false negative rate still decreases rapidly. The false negative rates of the uniformly and Student-t distributed data show a convergence similar to the one we observe for the Normal data. The Student-t data, however, exceeds the set confidence level for the false positive rate. A plausible explanation is that the heavy tails of the Student-t distribution produces, for small sample sizes, test statistics that are more extreme than what we expect under a normal noise assumption. For most of the experiments, however, we note that the false positive rate is quite below the selected confidence level. This is to be expected as the procedure can eliminate a false positive through multiple tests, leading to an overall conservative approach.
(a) Performance for $n=10$.
(b) Performance for $n=14$.
Figure 4: The effect of observing data in more environments on the performance strongly depends on the given sample size in each environment.
Increasing number of environments.
Next, we investigate the effect of increasing the number of environments. For that we use the SEM and Normal data and a sample size of $n=10$ and $n=14$. We report the results in Figure 4 and observe that for the small sample size of $n=10$ the performance on the SEM data is considerably worse than the performance on the Normal data. We note in particular that the false negative rate does not seem to converge for an increasing number of environments. This observation is not unique to the SEM data, and could also be observed for the Normal data for smaller sample sizes, as also predicted by Theorem 3. The gap in performance on the two datasets is mostly closed for a moderate increase of the sample size to $n=14$. The experiment shows that the importance of the number of available environments strongly depends on other factors of the given data, as for example the sample size per environment.
6.2 Comparison to Joint LiNGAM
We now consider a comparison of our method with joint LiNGAM (Shimizu, 2012). While the methods make different assumptions on the data, they are not exclusive. The main distinction is that joint LiNGAM assumes that the noise variables are non-normal distributed, while we assume that the target noise $\varepsilon^{e}$ has the same distribution for all $e\in[E]$. While joint LiNGAM can identify the complete causal graph, we generally can only detect the causal parents of $Y^{e}$. In our first experiment we highlight that our method can be very powerful in comparison to LiNGAM, even for small samples, while remaining conservative by controlling the false positive rate. We then show the caveat of our method, namely it will fail if the environments are not heterogeneous enough, something that leaves LiNGAM unaffected.
Increasing sample size.
We first compare the performance of LiNGAM and LoLICaP, with different sample sizes. We created data with the linear SEM from Appendix B with uniform (Uniform SEM) and Student-t (Student SEM) distributed noise. Figure 5 indicates that on the Uniform SEM data, LoLICaP quickly outperforms LiNGAM, while on the Student SEM data the reverse is the case.
(a) Performance comparison with data generated from the SEM with uniform noise.
(b) On the SEM data with Student-t noise LiNGAM outperforms LoLICaP.
Figure 5: The performance comparison between LoLICaP and joint LiNGAM, for different sample sizes. The left figure shows the performance on the data from the linear SEM with uniform noise and showcases the fast convergence of LoLICaP. The right shows the performance on the SEM data with Student-t noise, highlighting that the model-misspecification due to heavy tailed noise poses a problem for LoLICaP, in particular for the false positive rate.
(a) For $n=16$ samples LiNGAM and LoLICaP can make similar use of the additional environments.
(b) For $n=24$ LoLICaP can make more use of the added environments, eventually rendering the performance of the both methods comparable.
Figure 6: Comparison between LiNGAM and LoLICaP when both methods have access to additional environments. While in both experiments LiNGAM is able to make moderate improvements, LoLICaP can make especially good use of the additional environments for the larger sample size of $n=24$. This validates the findings discussed in Figure 4 and indicates that also under model-misspecifaction there is a complicated relationship between the performance and the number of available environments.
Increasing number of environments.
Next, we investigate the effect of adding more environments on the performance of LoLICaP And LiNGAM. We generated data from our SEM with Student-t noise with $n=16$ and $n=24$, since Figure 4(b) indicates that both LoLICaP and LiNGAM still can improve for those settings. We show the results for both cases in Figure 6. For the smaller sample size of $n=16$ both methods show a moderate improvement. Similarly to our observations in Figure 4, LoLICaP shows much better performance improvements with respect to the number of environments for the bigger sample size of $n=24$. While the performance of LiNGAM quickly saturates with an increased number of environments, LoLICaP can also benefit from larger number of environments eventually outperforming LiNGAM. A possible explanation for this in the context of this experiment: The parameters that are responsible for the heterogeneity across environments (the causal parameters and noise variances) are randomly sampled for each environment. Every new environment has then a chance to be sampled with parameters that strongly differ from all other environments, and in this sense each new environment increases the overall heterogeneity.
Reducing environment heterogeneity.
In the next experiment we highlight and emphasize the importance of the heterogeneity of the environments for our methodology. In this experiment the data was generated from the linear SEM with uniform noise and $n=20$, as described in Appendix B, but we change the heterogeneity of the variances of the covariates and the parameters $\beta^{e}$ along a heterogeneity parameter $h$. Here larger $h$ indicate stronger heterogeneity and $h=0$ indicates no heterogeneity across environments. We refer to Appendix B for a precise explanation of this parameter. As expected, and shown in Figure 6(a), the performance of LoLICaP quickly increases in $h$, while LiNGAM is unaffected by this parameter.
(a) Changing the heterogeneity of the environments with SEM data and uniform noise.
(b) Changing the heterogeneity of the environments with SEM data and Student-t noise.
Figure 7: Comparison between LiNGAM and LoLICaP when we change the heterogeneity of the environments. Here $h=0$ indicates that there was no heterogeneity across environments. While increasing the heterogeneity quickly helps LoLICaP to reach optimal performance in the uniform noise case (left), the performance for the Student-t noise is less affected (right).
6.3 Application: Network Detection in Dynamical Systems
Finally, we want to describe, and showcase, an application of LoLICaP that highlights its potential and current limitations: network detection in dynamical systems. Following the remarks after Theorem 1, finding a full causal graph is possible in our proposed setting if none of the covariates is directly affected by the environment index. In this section we simulate data from a non-linear dynamical system where this assumption indeed holds when we define our environments as different temporal intervals of length $n$. In this experiment we thus want to reveal if the non-linearity of the system can introduce sufficient heterogeneity across time so that LoLICaP can subsequently discover causal relations. More precisely, our data consists of a discrete-time and noisy version of a five-dimensional Lorenz system described by Shen (2014) together with an independently sampled random walk. The dynamical system is formally defined by the following set of equations, where the superscript $t$ indicates a time index.
$X^{t+1}_{1}=0.9X_{1}^{t}+0.1X^{t}_{2}+\varepsilon^{t}_{1}$ (10)
$X^{t+1}_{2}=0.28X_{1}^{t}-0.01X^{t}_{1}X^{t}_{3}+0.99X^{t}_{2}+\varepsilon^{t}_{2}$ (11)
$X^{t+1}_{3}=0.01X^{t}_{1}(X^{t}_{2}-X^{t}_{4})+0.9733X^{t}_{3}+\varepsilon^{t}_{3}$ (12)
$X^{t+1}_{4}=0.01X^{t}_{1}(X^{t}_{3}-2X^{t}_{5})+0.9366X_{4}^{t}+\varepsilon^{t}_{4}$ (13)
$X^{t+1}_{5}=0.02X^{t}_{1}X^{t}_{4}+0.96X^{t}_{5}+\varepsilon^{t}_{5}$ (14)
$X^{t+1}_{6}=X^{t}_{6}+\varepsilon^{t}_{6}$ (15)
Here $X^{t}_{1},\cdots,X^{t}_{5}$ defines the Lorenz system, while $X^{t}_{6}$ is the random walk. Furthermore $\varepsilon_{i}^{t}$ for $1\leq i\leq 6$ are random noise variables sampled independently from each other and past values from a standard normal distribution.
Given that we have dynamical data, we chose our environments to be time intervals of length $n$. More precisely, for a given starting time $t_{0}\in\mathbb{N}$ we define the observations in environment $e_{t_{0}}$ by $({X}^{e_{t_{0}}},{Y}^{e_{t_{0}}})\vcentcolon=\{(X^{t},Y^{t})\}_{t_{0}\leq t\leq t_{0}+n}$. The target variable $Y^{t}$ is now the observation of any chosen covariate, but at the next time-step. For example, if we want to find the causal parents of $X_{1}$, we define $Y^{t}\vcentcolon=X^{t+1}_{1}$.
(a) The reported causal graph when splitting the time series into intervals of length $n=20$. As the model is less misspecified than for $n=25$, we only report one two positives. This, however, comes at the cost of missing three connections.
(b) The reported causal graph when splitting the time series into intervals of length $n=25$. The model is more misspecified than for $n=20$, and with that we report three false positive causal parents. However, given the larger sample size we only miss one connection.
Figure 8: The causal graph reported by the procedure described in the text. Explanation of the arrows: Solid blue: Correctly found. Dashed red: Not found (false negative discovery). Dotted yellow: Falsely reported (false positive discovery.)
Experimental details.
First, over $500$ independent runs we generated $8500$ samples according to Equations (10)-(15), where the initial vector is set as
$(X^{0}_{1},X^{0}_{2},X^{0}_{3},X^{0}_{4},X^{0}_{5},X^{0}_{6})=(2,0.97,0.99,1,0.97,1).$
Given the data of one run, we want to split the time series into intervals of length $n$ and then run LoLICaP using those intervals as environments. For that, we need to decide on how to choose $n$. While one might envision data-driven choices, for these experiments we chose $n$ driven by the following rationale: If $n$ is chosen very small the algorithm tends to return the empty set because most subsets $S\subseteq[D]$ are plausible causal parents. This is in particular true if $n\leq 7$, so when the data can be interpolated. If $n$ is too large, the method tends also to return the empty set, as in that case no subset of covariates provides a set of plausible causal parents due to a strong violation of the local linearity assumption. We thus first tested for which sample sizes $n$ the method tends to not return the empty set for any covariate in an individual run. Setting $E=300$, this was generally the case for $n\in[15,35]$ which motivated our choice to set $n\in\{20,25\}$. Leaving the first $500$ samples as a warm-up phase for the system, we used the remaining samples to split the data into $E=300$ sequential environments of length $n\in\{20,25\}$. Over the $500$ runs we then count for each target how often each covariate was reported as a causal parent. In Appendix B.3 we report the complete counts over the $500$ runs of this experiment, for the sake of representation we report here a sparse causal network based on the following reasoning: If all assumptions of our model would be true, we expect by Proposition 2 that all covariates that LoLICaP (with $\alpha=0.1$) reported in over $10\%$ of the runs as causal parents are also true causal parents. We thus report all covariates as causal parents that were found significantly more often than the $10\%$ threshold. The significance is tested with a binomial test at significance level $0.05$, with the null that a covariate is found in $10\%$ of the runs, and the alternative that the reported rate is greater than $10\%$.
In Figure 8 we report the resulting causal graphs for $n\in\{20,25\}$ and we can affirm that non-linear relationships in a dynamical system can introduce sufficient heterogeneity across time to identify causal parents with LoLICaP. For $n=20$ we only report two false positive discoveries, while still being able to recover $14$ out of the $17$ true causal connections. While increasing the sample size to $n=25$ helps to recover all but one true causal connection, we also increase the false positive reports to $3$. By Proposition 2 we know that this has to be due to model misspecification, highlighting an important limitation of the local-linearity assumption, a problem that we further discuss in Section 7.
7 Discussion and Ongoing Work
In this paper we presented an extension of the work from Peters et al. (2016) to a setting with local linear models, with many interesting consequences. We now discuss limitations, how they can be addressed, and possible extensions of our work.
Scalability of the method.
One of the main computational issues of the proposed framework is that it scales exponentially with respect to the dimension $D$. There are several ways to deal with this issue. First, instead of looping over all possible subsets of $[D]$, as we do in Algorithm 3, one could greedily add or remove covariates as plausible causal parents. While the greedy removal was already applied to ICP by Salas-Porras et al. (2022), it is clear that such methodology can generally not enjoy the same guarantees as the full method. It is, however, not clear how these greedy methods perform in practice against the full method, and this poses an interesting and important open question. Second, one can reduce the dimensionality $[D]$, for example by clustering highly correlated variables. Instead of looping over all subsets from $[D]$, one may then loop over all clusters. While this reduces the dimensionality, one then does not find the individual causal parents, but instead the clusters in which a parent is present. Next, if it is sufficient to not find exclusively the direct causal parents, but also other ancestors, Mogensen et al. (2022) provide a scalable alternative by only considering minimal sets of plausible parents. Finally, we envision a procedure where the test-criterion given by $\phi_{S}$ is encoded as an objective function that one may optimize over the set of covariates, in the spirit of procedures such as the LASSO. While we are not aware of any work that does this in the context we consider, similar ideas have been applied in the machine learning literature, see for example Arjovsky et al. (2019).
The role of locality.
The main novelty of our proposed setting is that we model each environment separately with a local model without any additional structural assumptions between the local models. This relaxes the global linearity assumption and, in some sense, allows us to model non-linear systems as seen in Section 6. The linearity relaxation, however, is not the only strength of this setting: In some scenarios the data generation in different environments can truly follow different functional relationships, and in that case no global model, linear or non-linear, exists. If we, for example, define different car types as different environments, one can assume that there is not one fixed model that can explain the behavior of the parts of all those cars, while only the causal relationships between the parts are the same. It then becomes necessary to model the environments separately, no matter how expressive the model class is.
Limitations of local linearity.
In the experiments of Section 6.3 we approximated a non-linear dynamical systems with intervals of linear models. To ensure approximate linearity we had to set the length of the intervals, which corresponds to the sample size $n$ of the environments within our setting, at a relatively small number of $n=20$. In that experiment this sample size was too small to recover some variables as causal parents, while some false discoveries are likely due the violation of the linearity assumption. Both problems highlight the need for non-linear versions of the proposed methodology. While there are efforts towards more general models, as also investigated by Heinze-Deml et al. (2018), the proposed framework offers a path to use different kind of machine learning methods to model the causal relationships. While the specific methodology proposed in this paper will no longer apply, as we do not know the behavior of the residuals for general machine learning methods, the conceptual idea still makes sense: If, for a given set $X^{e}_{S}$ of covariates, we would have access to an oracle optimal model $Y^{e}=f_{S}(X^{e}_{S})$, one still expects that some properties of $(Y^{e},f_{S}(X^{e}_{S}))$ are generally invariant across $e\in[E]$ only if $S^{*}\subseteq S$. The main challenges here are to estimate $f_{S}$, which can be very problematic if $f_{S}$ belongs to a complex function class, and to find meaningful properties that we can test for invariance, like the residuals for the linear case.
Changing distribution of the target noise.
Arguably one of the strongest assumptions in our setting is that the target noise $\varepsilon^{e}$ follows the same distribution in every environment. While this ensures that the environment cannot have any influence on $\varepsilon^{e}$, this assumption can be relaxed. We may, for example, parameterize the distribution of $\varepsilon^{e}$ along a parameter $\theta^{e}$ and then assume that $\theta^{e}$ and the environment indices are independently sampled from a distribution. With that assumption, the environment index is independent of $\theta^{e}$ and thus should not have any effect on the distribution of $\varepsilon^{e}$. Establishing that a set $S\subseteq[D]$ is plausible then can, for example, be established by testing the independence of the environment index and $\theta^{e}$.
Picking good environments.
In applications we may often face the choice of how to define the different environments in which our data is partitioned. In the experiment of Section 6.3, for example, we have access to a stream of dynamical data and need to split this stream in a meaningful way. While in this experiment we simply picked environments that are evenly spaced in time, one can think about more sophisticated ways of picking environments: We know that a set of good environment fulfills approximately the local linearity assumption and are heterogeneous from each other. We believe that those assumptions are, at least partially, testable, and identifying good environments should be a focus of future research. This is not only important for our own methodology, but any method that views an environment as an accidental interventions.
Acknowledgements
This research has been funded by NWO under the grant PrimaVera (https://primavera-project.com) number NWA.1160.18.238 and was done in collaboration with our project partner ASML.
Appendix A Proofs of main results
A.1 Proof of Lemma 1
Lemma 1
If $\tilde{H}_{0,S}$ is true, then so is $H_{0,S}$.
Proof.
If $\tilde{H}_{0,S}$ is true then, using the independence of the residual noise $r^{e}$ and covariates, we know that $\gamma^{e}$ as well as $\tilde{\beta}^{e}_{S}$ are solutions of the least squares problem
$\min\limits_{\beta}\mathbf{E}\left[\sum_{i=1}^{n^{e}}(X^{e}_{S,i}\beta-Y^{e}_{i})^{2}\right].$
This means that $X^{e}_{S,i}\tilde{\beta}=X^{e}_{S,i}\gamma^{e}$ almost surely, which implies the lemma. ∎
A.2 Proof of Theorem 2
We first collect some lemmas that are needed for the main proof.
Lemma 2.
For any two random variables $X,Y$ and $c>0$ it holds that
$\mathbf{P}(X<Y)\leq 2(\mathbf{P}(X<c)+\mathbf{P}(Y>c)).$
Proof.
First we define the three events
$A=\{(X<c\land Y<c)\land(X<Y)\}$
$B=\{(X>c\land Y>c)\land(X<Y))\}$
$C=\{X<c\land Y>c\}.$
With that we may conclude that
$\mathbf{P}(X<Y)$ $=\mathbf{P}(A\lor B\lor C)=\mathbf{P}(A)+\mathbf{P}(B)+\mathbf{P}(C)$
$\leq\mathbf{P}(X<c)+\mathbf{P}(B)+\mathbf{P}(C)$
$\leq\mathbf{P}(X<c)+\mathbf{P}(Y>c)+\mathbf{P}(X<c\land Y>c)$
$=\mathbf{P}(X<c)+\mathbf{P}(Y>c)+\mathbf{P}(X<c)+\mathbf{P}(Y>c)-\mathbf{P}(X<c\lor Y>c)$
$\leq 2(\mathbf{P}(X<c)+\mathbf{P}(Y>c)).$
Lemma 3 (Adapted from Dasgupta and Gupta (2003), Lemma 2.2).
Let $Z\sim\chi^{2}(k)$ and $F_{Z}$ be the cumulative distribution function of $Z$. Then then following two inequalities hold:
$1-F_{Z}(kz)\leq(z)^{\frac{k}{2}}e^{\frac{k}{2}(1-z)}\text{\ for $z>1$ }$ (16)
$F_{Z}(kz)\leq(z)^{\frac{k}{2}}e^{\frac{k}{2}(1-z)}\text{\ for $0<z<1$ }.$ (17)
Proof.
We begin with the second inequality. For $1\leq i\leq k$ let $X_{i}\sim\mathcal{N}(0,1)$ be $k$ independent standard normal random variables. Then $Z\vcentcolon=\sum^{k}_{i=1}X_{i}^{2}\sim\chi^{2}(k)$. We use a Chernoff bounding technique as follows and for $t>0$ we derive
$F_{Z}(kz)$ $=\mathbf{P}(kz-\sum^{k}_{i=1}X_{i}^{2}\geq 0)=\mathbf{P}(e^{tkz-t\sum^{k}_{i=1}X_{i}^{2}}\geq 1)$
$\leq e^{tkz}\mathbf{E}\left[e^{-t\sum^{k}_{i=1}X_{i}^{2}}\right]=e^{tkz}\mathbf{E}\left[e^{-tX_{i}^{2}}\right]^{k}.$
Using the known equality $\mathbf{E}\left[e^{-tX_{i}^{2}}\right]=(1+2t)^{-\frac{1}{2}}$ for $-\frac{1}{2}<t<\infty$ we may set $t=\frac{1}{2}\frac{1-z}{z}$ since $z<1$ to obtain
$e^{tkz}\mathbf{E}\left[e^{-tX_{i}^{2}}\right]^{k}=e^{tkz}(1+2t)^{-\frac{k}{2}}=e^{\frac{k}{2}(1-z)}z^{\frac{k}{2}}.$
The first inequality of the lemma follows similarly. ∎
Lemma 4.
Let $S\subset[D]$ with $S^{*}\not\subseteq S$ and set $U\vcentcolon=S\setminus S^{*}$. Then, under Assumptions 1, 3 and 4 we have the following properties of the SSR when regressing ${Y}$ onto ${X}_{S}$ in environments $v\in[E_{1}]$ and $w\in[E_{2}]$. Defining $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}$ it holds that $\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$. Here $r^{v}_{S}$ and $r^{w}_{S}$ are the residuals defined in Algorithm 3.
Proof.
Regressing the target ${Y}$ only on the covariates in $S$ we obtain in environment $v\in[E_{1}]$ the linear model
${Y}^{v}=\beta^{1}_{S}{X^{v}_{S}}+r^{v}_{S},$
where $r^{v}_{S}=\beta^{1}_{U}{X^{v}_{U}}+\varepsilon^{e}$. Because of the normality and independence assumptions, we find that $r^{v}_{S}$ follows a normal distribution with variance $\rho^{v}=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$. Furthermore $r^{v}_{S}$ can be assumed to be of zero mean, due to our inclusion of a column of constant ones in $X^{v}_{S}$. This implies that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(k)$ and we can define the equivalent expression for $w\in[E_{2}]$ by setting $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(k)$. ∎
Proof of Theorem 2.
Proof.
To identify when our method accepts $S$ as a set of potential causal parents, we need to control the probability $\mathbf{P}(\phi_{S}=0)$. For convenience, we introduce $Z_{\min}\vcentcolon=\min_{e\in[E]}Z^{e}_{S}$ and $Z_{\max}\vcentcolon=\max_{e\in[E]}Z^{e}_{S}$. The probability of a false negative can be bounded by Markov’s Inequality as
$\mathbf{P}(\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0)$ $=\mathbf{P}\left(\mathbf{P}\left(T(\boldsymbol{X},\boldsymbol{Y})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\ \bigg{\lvert}\ \boldsymbol{X},\boldsymbol{Y}\right)\geq\alpha\right)$ (18)
$\leq\frac{1}{\alpha}\mathbf{P}\left(T(\boldsymbol{X},\boldsymbol{Y})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right).$ (19)
We continue with analyzing the quantity $\mathbf{P}(T(\boldsymbol{X},\boldsymbol{Y})>{Z_{\text{min}}}/{Z_{\text{max}}})$ and in particular try to understand the distribution of $T(\boldsymbol{X},\boldsymbol{Y})$ under Assumption 4. By Lemma 4 we know that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ for $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}$. This allows us to write:
$T(\boldsymbol{X},\boldsymbol{Y})=\frac{\min_{v\in[E_{1}],w\in[E_{2}]}(\min(\rho^{v}{P}^{v},\rho^{w}{Q}^{w}))}{\max_{v\in[E_{1}],w\in[E_{2}]}(\max(\rho^{v}{P}^{v},\rho^{w}{Q}^{w}))}.$
With the help of Lemma 2 we may then for any $c>0$ bound
$\mathbf{P}\left(T(\boldsymbol{X},\boldsymbol{Y})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)$ $\leq\mathbf{P}\left(\frac{\min\limits_{w\in[E_{2}]}\rho^{w}{Q}^{w}}{\max\limits_{v\in[E_{1}]}\rho^{v}{P}^{v}}>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)$
$\leq 2\mathbf{P}\left(\frac{\min\limits_{w\in[E_{2}]}\rho^{w}{Q}^{w}}{\max\limits_{v\in[E_{1}]}\rho^{v}{P}^{v}}>c\right)+2\mathbf{P}\left(c>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right).$ (20)
With the above inequality we are allowed to continue with the expression in line (20) and we start with the first term $\mathbf{P}({\min\limits_{w\in[E_{2}]}\rho^{w}{Q}^{w}}/{\max\limits_{v\in[E_{1}]}\rho^{v}{P}^{v}}>c)$. By setting $c=\sqrt{\frac{\rho^{w}}{\rho^{v}}}<1$ we have that $c\frac{\rho^{v}}{\rho^{w}}=\frac{1}{c}$, which allows us to write
$\mathbf{P}\left(\frac{\min\limits_{w\in[E_{2}]}{Q}^{w}}{\max\limits_{v\in[E_{1}]}{P}^{v}}>c\frac{\rho^{v}}{\rho^{w}}\right)=\mathbf{P}\left(\frac{\max\limits_{v\in[E_{1}]}{P}^{v}}{\min\limits_{w\in[E_{2}]}{Q}^{w}}<c\right),$
which is now the quantity we study further. We further want to simplify this term by splitting it with the help of the two events
$\mathcal{E}_{1}\vcentcolon=\left\{\frac{\max\limits_{v\in[E_{1}]}{P}^{v}}{\min\limits_{w\in[E_{2}]}{Q}^{w}}<c\right\}$
$\mathcal{E}_{2}\vcentcolon=\left\{{\max\limits_{v\in[E_{1}]}{P}^{v}}>k\sqrt{c}\ \land\ \min\limits_{w\in[E_{2}]}{Q}^{w}<k\frac{1}{\sqrt{c}}\right\}.$
Noting that $\mathbf{P}(\mathcal{E}_{1},\mathcal{E}_{2})=0$ we can bound
$\mathbf{P}\left(\frac{\min\limits_{w\in[E_{2}]}{Q}^{w}}{\max\limits_{v\in[E_{1}]}{P}^{v}}>c\frac{\rho^{v}}{\rho^{w}}\right)$ $=\mathbf{P}\left(\frac{\max\limits_{v\in[E_{1}]}{P}^{v}}{\min\limits_{w\in[E_{2}]}{Q}^{w}}<c\right)=\mathbf{P}(\mathcal{E}_{1})$
$\leq\mathbf{P}(\mathcal{E}_{1},\mathcal{E}_{2})+\mathbf{P}(\{\mathcal{E}_{2}\}^{c})$ (21)
$\leq\mathbf{P}\left({\max\limits_{v\in[E_{1}]}{P}^{v}}<k\sqrt{c}\right)+\mathbf{P}\left({\min\limits_{w\in[E_{2}]}{Q}^{w}}>k\frac{1}{\sqrt{c}}\right).$ (22)
With the reminder that $c<1$ and that for any $v\in[E_{1}]$ and $w\in[E_{2}]$ the terms $P^{v}$ and $Q^{w}$ follow a Chi-square distribution with $k$ degrees of freedom we can use Lemma 3 to conclude that for any $v_{0}\in[E_{1}]$
$\mathbf{P}\left({\max\limits_{v\in[E_{1}]}{P}^{v}}<k\sqrt{c}\right)\leq\mathbf{P}(P^{v_{0}}<k\sqrt{c})\leq(\sqrt{c})^{\frac{k}{2}}e^{\frac{k}{2}(1-\sqrt{c})}.$ (23)
And similarly, we derive
$\mathbf{P}\left({\min\limits_{w\in[E_{2}]}{Q}^{w}}>k\frac{1}{\sqrt{c}}\right)\leq\left(\frac{1}{\sqrt{c}}\right)^{\frac{k}{2}}e^{\frac{k}{2}\left(1-\frac{1}{\sqrt{c}}\right)}.$ (24)
With this we can control the first term of our intermediate target defined in (20) as plugging Inequalities (23) and (24) into (22) provides the result
$\mathbf{P}\left(\frac{\max\limits_{v\in[E_{1}]}{P}^{v}}{\min\limits_{w\in[E_{2}]}{Q}^{w}}<c\right)\leq(\sqrt{c})^{\frac{k}{2}}e^{\frac{k}{2}(1-\sqrt{c})}+\left(\frac{1}{\sqrt{c}}\right)^{\frac{k}{2}}e^{\frac{k}{2}\left(1-\frac{1}{\sqrt{c}}\right)}.$ (25)
The other term in our target (20) is given by $\mathbf{P}({Z_{\text{min}}}/{Z_{\text{max}}}<c)$. To bound this term, we can use almost the exact same reasoning as for the first term, the only difference being the dependence on $E$ as for this term we use a union bound in the inequalities that correspond to (23) and (24) for the previous term. In the end we obtain that
$\mathbf{P}\left(\frac{Z_{\text{min}}}{Z_{\text{max}}}<c\right)\leq E\left((\sqrt{c})^{\frac{k}{2}}e^{\frac{k}{2}(1-\sqrt{c})}+\left(\frac{1}{\sqrt{c}}\right)^{\frac{k}{2}}e^{\frac{k}{2}\left(1-\frac{1}{\sqrt{c}}\right)}\right).$ (26)
Plugging the result from (25) and (26) back into (20) we obtain
$\mathbf{P}\left(T(\boldsymbol{X},\boldsymbol{Y})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)\leq 4E\left((\sqrt{c})^{\frac{k}{2}}e^{\frac{k}{2}(1-\sqrt{c})}+\left(\frac{1}{\sqrt{c}}\right)^{\frac{k}{2}}e^{\frac{k}{2}\left(1-\frac{1}{\sqrt{c}}\right)}\right).$
Finally, plugging this back into Inequality (19) and noting that $c=\sqrt{\frac{\rho^{w}}{\rho^{v}}}$, we obtain the statement of the theorem. ∎
A.3 Proof of Theorem 3
First, we state two useful lemmas needed for the proof. We do not claim originality on those statements as those type of derivations may be found in literature on extremal events such as Embrechts et al. (2013). As we could, however, not find the precise statements needed, we prove them now.
Lemma 5.
Let $q\in\mathbb{N}$ and for $e\in[E]$ let $C_{1}^{e}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\chi^{2}(k)$ and for $e\in[qE]$ let $C_{2}^{e}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\chi^{2}(k)$. For the random variables $Q_{E}=\max\limits_{e\in[E]}C_{1}^{e}$ and $W_{E}=\max\limits_{e\in[qE]}C^{e}_{2}$ it then holds that
$\lim\limits_{E\to\infty}\mathbf{P}\left(\frac{W_{E}}{Q_{E}}=1\right)=1.$
Proof.
We prove the lemma by showing that for any $c>1$ we have $\lim\limits_{E\to\infty}\mathbf{P}({W_{E}}/{Q_{E}}>c)=0$ and for any $c<1$ that $\lim\limits_{E\to\infty}\mathbf{P}({W_{E}}/{Q_{E}}<c)=0$. The statement of the lemma then follows from a union bound over the events $\{{W_{E}}/{Q_{E}}\not\in[1-\frac{1}{m},1+\frac{1}{m}\}_{m\in\mathbb{N}}$. Hashorva et al. (2012) show that there exists a series $b_{E}$ such that for $E\to\infty$ we have $b_{E}\to\infty$ and $(W_{E}-b_{E})$ converges to a distribution with support on $\mathbb{R}$. With that in hand we start by showing the case for $c>1$. First, choose $\delta>0$ such that $c>\frac{1+\delta}{1-\delta}$. With this we have that
$\mathbf{P}\left(\frac{W_{E}}{Q_{E}}>c\right)$ $\leq\mathbf{P}\left(\frac{W_{E}}{Q_{E}}>\frac{b_{E}(1+\delta)}{b_{E}(1-\delta)}\right)$
$\leq\mathbf{P}(W_{E}>b_{E}(1+\delta))+\mathbf{P}(Q_{E}<b_{E}(1-\delta)).$
We observe that for $E\to\infty$ the probability $\mathbf{P}(W_{E}>b_{E}(1+\delta))=\mathbf{P}((W_{E}-b_{E})>b_{E}\delta)$ converges to 00 since $(W-b_{E})$ converges to a distribution with support on $\mathbb{R}$ and $b_{E}\to\infty$. Analogue to this one may show that $\lim\limits_{E\to\infty}\mathbf{P}(Q_{E}<b_{E}(1-\delta))=0$. The case for $c<1$ works analogue to $c>1$. ∎
Lemma 6.
For $e\in[E]$ let $W^{e}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\chi^{2}(k)$. Then
$\lim\limits_{E\to\infty}\mathbf{P}\left(E^{\frac{2}{k}}\min\limits_{e\in[E]}W^{e}>w\right)=e^{-w^{\frac{k}{2}}c_{0}},$
where $c_{0}\in\mathbb{R}$ is a term constant in $w$. This also implies that the density function $f_{E}(w)$ of the random variable $\lim\limits_{E\to\infty}E^{\frac{2}{k}}\min\limits_{e\in[E]}W^{e}$ is given by
$f(w)=w^{\frac{k}{2}-1}c_{0}e^{-w^{\frac{k}{2}}c_{0}}.$
Proof.
The cumulative distribution function $F(w)$ of any $W^{e}$ is given by $F(w)=\tilde{c}_{0}\gamma\left(\frac{k}{2},\frac{w}{2}\right),$ where $\tilde{c}_{0}\vcentcolon=\frac{1}{\Gamma(\frac{k}{2})}$ is a term constant in $w$, Here $\Gamma$ is the gamma function and $\gamma$ is the lower incomplete gamma function defined for $s>0,x>0$ as
[…]
Using the power series definition of the exponential term we can derive that
$\gamma\left(\frac{k}{2},\frac{w}{2}\right)=\left(\frac{w}{2}\right)^{\frac{k}{2}}\sum_{m\geq 0}\left(\frac{-w}{2}\right)^{m}\frac{1}{m!(\frac{k}{2}+m)}=\frac{2}{k}\left(\frac{w}{2}\right)^{\frac{k}{2}}+O\left(w^{\frac{k}{2}+1}\right).$
Next, using the relation $F(w)=\tilde{c}_{0}\gamma\left(\frac{k}{2},\frac{w}{2}\right)$ and the equation above we conclude that for $c_{0}\vcentcolon=\frac{2}{k}\tilde{c}_{0}$ it holds that
$\lim\limits_{E\to\infty}\mathbf{P}\left(E^{\frac{2}{k}}\min\limits_{e\in[E]}W^{e}>w\right)=\lim\limits_{E\to\infty}\left(1-F\left(\frac{w}{E^{\frac{2}{k}}}\right)\right)^{E}$
$=\lim\limits_{E\to\infty}\left(1-\frac{1}{E}c_{0}w^{\frac{k}{2}}-O\left(E^{-\left(1+\frac{2}{k}\right)}\right)\right)^{E}=e^{-c_{0}w^{\frac{k}{2}}}.$
The last equality uses the limit definition of the exponential function.
Proof of Theorem 3
Proof.
To identify when our method accepts $S$ as a set of potential causal parents, we have to understand the probability $\mathbf{P}(T_{S}>{\min_{e\in[E]}Z^{e}_{S}}/{\max_{e\in[E]}Z^{e}_{S}}$), which is the key quantity for the hypothesis test $\phi_{S}$, defined in Equation (9), that our Algorithm 3 uses. For notational convenience we set $Z_{\text{min}}\vcentcolon=\min\limits_{e\in[E]}Z^{e}$ and $Z_{\text{max}}\vcentcolon=\max\limits_{e\in[E]}Z^{e}$. For this proof we assume without loss of generality that $\rho^{v}>\rho^{w}$. We are investigating the behavior of $\mathbf{P}(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>{Z_{\text{min}}}/{Z_{\text{max}}})$ for $E\to\infty$, the statements of the theorem will then follow by Markov’s Inequality. By Lemma 4 we know that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ for $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}$. This allows us to rewrite $T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})$ as:
$T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})=\frac{\min_{v\in[E_{1}],w\in[E_{2}]}(\min(\rho^{v}{P}^{v},\rho^{w}{Q}^{w}))}{\max_{v\in[E_{1}],w\in[E_{2}]}(\max(\rho^{v}{P}^{v},\rho^{w}{Q}^{w}))}.$
To proof the first statement of the theorem we note that
$\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)$ $\leq\mathbf{P}\left(\frac{\min_{w\in[E_{2}]}\rho^{w}Q^{w}}{\max_{v\in[E_{1}]}\rho^{v}P^{v}}>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)$
$=\mathbf{P}\left(\frac{\rho^{w}}{\rho^{v}}\frac{\min_{w\in[E_{2}]}Q^{w}}{Z_{\text{min}}}>\frac{\max_{v\in[E_{1}]}P^{v}}{Z_{\text{max}}}\right).$
By Lemma 5 it holds that
$\lim\limits_{E\to\infty}\mathbf{P}\left(\frac{\rho^{w}}{\rho^{v}}\frac{\min_{w\in[E_{2}]}Q^{w}}{Z_{\text{min}}}>\frac{\max_{v\in[E_{1}]}P^{v}}{Z_{\text{max}}}\right)=\lim\limits_{E\to\infty}\mathbf{P}\left(\frac{\rho^{w}}{\rho^{v}}\frac{\min_{w\in[E_{2}]}Q^{w}}{Z_{\text{min}}}>1\right).$ (27)
For brevity we write $c\vcentcolon=\frac{\rho^{v}}{{\rho^{w}}}=\frac{1}{I_{S}}$ and $Q\vcentcolon=\min_{w\in[E_{2}]}Q^{w}$. Defining $f_{E}(x)$ as the density function of $E^{\frac{2}{k}}Z_{\text{min}}$ and using the result of Lemma 6 we continue with:
$\lim\limits_{E\to\infty}\mathbf{P}\left(\frac{Q}{Z_{\text{min}}}>c\right)$ $=\lim\limits_{E\to\infty}\mathbf{P}\left(\frac{E^{\frac{2}{k}}Q}{E^{\frac{2}{k}}Z_{\text{min}}}>c\right)$
[…]
$=-\frac{2}{c^{\frac{k}{2}}+2}e^{-x^{\frac{k}{2}}c_{0}\left(\frac{1}{2}c^{\frac{k}{2}}+1\right)}\biggr{\rvert}^{\infty}_{0}=\frac{2}{c^{\frac{k}{2}}+2}=\frac{2(I_{S})^{\frac{k}{2}}}{2(I_{S})^{\frac{k}{2}}+1}.$
Here we use from (28) to (29) Lebesgue’s dominated convergence theorem, which allows us to move the limit into the integral, and the limiting results from Lemma 6. To summarize, we have shown that
$\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)\leq\frac{2(I_{S})^{\frac{k}{2}}}{2(I_{S})^{\frac{k}{2}}+1}.$
The first statement of the theorem then follows by using this bound, together with the Markov’s Inequality applied to
$\mathbf{P}(\phi_{S}(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})=0)=\mathbf{P}\left(\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\ \bigg{|}\ \boldsymbol{X}_{E},\boldsymbol{Y}_{E}\right)\geq\alpha\right).$
To proof the second statement of the theorem we first note that
$\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)$
$\geq\mathbf{P}\left(\frac{\min_{v\in[E_{1}],w\in[E_{2}]}(\rho^{w}\min({P}^{v},{Q}^{w}))}{\max_{v\in[E_{1}],w\in[E_{2}]}(\rho^{v}\max({P}^{v},{Q}^{w}))}>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right),$
making use of the assumption that $\rho^{v}>\rho^{w}$. The normality and independence Assumption 3 together with the additional mutual independence assumption of the collection $\{P^{v},Q^{w}\}_{v\in[E_{1}],w\in[E_{2}]}$ allows us again to apply Lemma 5 and similar derivations to the ones following Equation (27) then lead to the conclusion that
$\lim\limits_{E\to\infty}\mathbf{P}\left(\frac{\min_{v\in[E_{1}],w\in[E_{2}]}(\rho^{w}\min({P}^{v},{Q}^{w}))}{\max_{v\in[E_{1}],w\in[E_{2}]}(\rho^{v}\max({P}^{v},{Q}^{w}))}>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)=\frac{(I_{S})^{\frac{k}{2}}}{(I_{S})^{\frac{k}{2}}+1}.$
In summary this means that
$\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\right)\geq\frac{(I_{S})^{\frac{k}{2}}}{(I_{S})^{\frac{k}{2}}+1}.$
The second statement of the theorem follows if we combine the bound above together with a transformation of the bound given by the following Markov’s Inequality:
$1-\mathbf{P}\left(\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\ \bigg{|}\ \boldsymbol{X}_{E},\boldsymbol{Y}_{E}\right)\geq\alpha\right)$
$=\mathbf{P}\left(1-\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\ \bigg{|}\ \boldsymbol{X}_{E},\boldsymbol{Y}_{E}\right)\geq 1-\alpha\right)$
$\leq\frac{1}{1-\alpha}\left(1-\mathbf{E}\left[\mathbf{P}\left(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>\frac{Z_{\text{min}}}{Z_{\text{max}}}\ \bigg{|}\ \boldsymbol{X}_{E},\boldsymbol{Y}_{E}\right)\right]\right).$
Appendix B Data Generation and Additional Results
In this section we describe the precise data generation mechanisms used in our experiments. Unless otherwise stated we fixed $E=30$, $D=6$ and $|S^{*}|=2$. In Section 6.1 and 6.2 we also experiment with data that was generated with a linear structural equation model (SEM). The SEM is defined by Equations (30)-(36) further below, with $\mathcal{D}(\sigma)$ being a distribution with zero mean and standard deviation $\sigma$. The specific choices of $\sigma$ for the individual experiments are specified in the following subsections. The graphical representation of the SEM is shown in Figure 9.
$X_{1}=\mathcal{D}(0,\sigma_{1})$ (30)
$X_{2}=X_{1}+\mathcal{D}(0,\sigma_{2})$ (31)
$X_{3}=0.3X_{1}+\mathcal{D}(0,\sigma_{3})$ (32)
$X_{4}=0.2X_{3}+\mathcal{D}(0,\sigma_{3})$ (33)
$Y=\beta_{2}X_{2}+\beta_{3}X_{3}+\mathcal{D}(0,\sigma_{Y})$ (34)
$X_{5}=0.1X_{2}+Y+\mathcal{D}(0,\sigma_{5})$ (35)
$X_{6}=Y+\mathcal{D}(\sigma_{6})$ (36)
B.1 Data from Section 6.1
The data for a single run is generated in the following way. For each environment $e\in[E]$ we sample a vector of standard deviations $(\sigma^{e}_{1},\cdots,\sigma^{e}_{6})$, such that each entry is independently sampled from the uniform distribution on $[1,5]$. The support entries of $\beta^{e}$, which are given by $S^{*}=\{2,3\}$, are sampled in the same way. Similarly, a vector of means $(m_{1}^{e},\cdots,m_{6}^{e})$ is entry-wise sampled independently from the uniform distribution on $[-1,1]$. The entries of ${X}^{e}_{d}$ are then an i.i.d $n$-sample of $\mathcal{D}(m^{e}_{d},\sigma^{e}_{d})$, where the specific distribution is specified in the experiments, i.e., $\mathcal{D}(m^{e}_{d},\sigma^{e}_{d})$ is either a normal, uniform or student-t distribution with mean $m^{e}_{d}$ and standard deviation $\sigma^{e}_{d}$. The target is defined as ${Y}^{e}=\beta^{e}{X}^{e}+{\varepsilon}^{e}$, such that ${\varepsilon}^{e}$ is an i.i.d $n$-sample from $\mathcal{D}(0,2)$. The data of SEM is created similarly, but with the additional relations defined through Equation (30)-(36).
B.2 Data from Section 6.2
[…]
Figure 9: The structure of the linear structural equation model (SEM) we use in some experiments, ignoring the noise variables. The corresponding structural equations are given in (30)-(36).
B.3 Full Results of the Network Detection
Here we report the full counts of the experiments of Section 6.3. More precisely, let $\tilde{S}_{r,j}$ be the set of causal parents that LoLICaP with given sample size $n$ reported in run $r$ for target covariate $j$, then we define $M^{n}_{i,j}:=\sum_{r=1}^{500}\mathbf{1}\left\{i\in\tilde{S}_{r,j}\right\}$. The results of the experiments from Section 6.3 are then given by the following two matrices
$M^{20}=\begin{bmatrix}494&23&7&6&9&3\\ 43&489&130&13&19&17\\ 54&108&442&371&44&24\\ 109&90&120&362&361&42\\ 82&81&53&405&405&41\\ 4&8&3&8&6&490\end{bmatrix},M^{25}=\begin{bmatrix}498&56&3&3&3&3\\ 68&470&238&28&27&17\\ 87&193&337&320&97&43\\ 110&93&108&189&189&44\\ 96&72&58&227&227&38\\ 2&5&3&6&3&496\end{bmatrix}.$